### v1

# PAE MobiLLM: Privacy-Aware and Efficient LLM Fine-Tuning on the Mobile Device via Additive Side-Tuning

Abstract.
There is a huge gap between numerous intriguing applications fostered by on-device large language model (LLM) fine-tuning (FT) from fresh mobile data and the limited resources of a mobile device. While existing server-assisted methods (e.g., split learning or side-tuning) may enable LLM FT on the local mobile device, they suffer from heavy communication burdens of activation transmissions, and may disclose data, labels or fine-tuned models to the server. To address those issues, we develop PAE MobiLLM, a privacy-aware and efficient LLM FT method which can be deployed on the mobile device via server-assisted additive side-tuning. The proposed PAE MobiLLM is featured by (i) notable communication and computing efficiency improvement; (ii) keeping data, true labels, and fine-tuned full models on the local mobile device; (iii) fast side-tuning on the server side. Briefly, PAE MobiLLM divides LLM fine-tune workloads between the mobile device and the server, i.e., the mobile device merely retains a frozen backbone model, while the computation-intensive trainable side-network is offloaded to the server. To further accelerate FT convergence and improve computing efficiency, PAE MobiLLM integrates activation caching on the server side, which allows the server to reuse historical activations and saves the mobile device from repeatedly computing forward passes for the recurring data samples. Besides, to reduce communication cost, PAE MobiLLM develops a one-token (i.e., “pivot” token) activation shortcut that transmits only a single activation dimension instead of full activation matrices to guide the side network tuning. Last but not least, PAE MobiLLM introduces the additive adapter side-network design which makes the server train the adapter modules based on device-defined prediction differences rather than raw ground-truth labels. In this way, the server can only assist device-defined side-network computing, and learn nothing about data, labels or fine-tuned models. Extensive experimental results demonstrate PAE MobiLLM’s superiority.
LLM fine-tuning, Mobile device, Side-tuning, Computing and communication efficiency, Privacy awareness.
Both authors contributed equally to this research.
ccs: Computing methodologies Distributed artificial intelligence
1. Introduction
Large language models (LLMs) such as GPT (2), RoBERTa (17) and LLaMA (27) have recently demonstrated remarkable capabilities across various NLP tasks, including language representation (22), cross-task reasoning (34), and context-aware generation (18). This success stems from their two-stage adaptation paradigm: pretraining on large-scale public corpora followed by domain-specific fine-tuning (FT). While effective in centralized cloud environments, this conventional approach faces fundamental challenges in mobile computing scenarios, where user data is continuously generated across distributed mobile devices (e.g., smartphones and wearables), highly privacy-sensitive (e.g., health monitoring records and personal chat histories, and subject to stringent latency demands for intelligent services (15; 33). By adapting a local LLM with fresh data on the mobile device, on-device fine-tuning is emerging as both a practical necessity and a strategic shift toward personalized intelligence without compromising user privacy.
Despite those promising applications, on-device LLM FT faces significant computing resource challenges. Because LLMs scale to billions of parameters, both the computational and memory requirements for fine-tuning an LLM can be prohibitively high for existing mobile devices’ hardware. Beyond storing the full model parameters, the predominant LLM training algorithm necessitates substantial memory for intermediate activations, gradients, and optimizer states. For example, fine-tuning a 7B-parameter LLM typically requires over 70 GB of GPU memory, far exceeding the capacity of most mobile devices, which often have 4–12 GB of RAM (32). A single training epoch for such an LLM could require hours of processing on mobile GPUs (3), which means full on-device fine-tuning cycles will take multiple days. This memory usage and latency fundamentally conflict with practical deployment scenarios requiring continuous model adaptation.
To enable LLM FT on a computing resource constrained mobile device, researchers are exploring to get some help from the high-performance computing server and collaborate the mobile device and the server via wireless communications supported by current mobile edge computing infrastructures. An intuitive solution is to integrate split learning (11) with PEFT, i.e., training shallow layers on the mobile device while offloading deeper layers to the server as shown in Fig. 1(a). Yet, the sequential processing and bidirectional communications in this approach require fast wireless access and strict synchronization. Another device-server collaboration LLM FT approach is MobiLLM (16), which introduces server-assisted parallel side-tuning. Instead of splitting the model vertically by layers, MobiLLM decouples the trainable modules (hosted on the server) from the frozen backbone model (retained on the mobile device) as shown in Fig. 1(b). During LLM FT, the mobile device executes forward passes on the backbone model and transmits intermediate activations to the server. The server computes gradients and updates the side network via backpropagation, which helps the mobile device bypass the computing and memory intensive gradient calculations. By eliminating backpropagation, MobiLLM enables the LLM FT of billion-parameter models (e.g., OPT-1.3B) on the mobile device within 4.5 GB memory.
Although the pioneering device-server collaboration framework relieves the memory concerns of mobile devices, it incurs new challenges for the practical deployment of LLM FT on the mobile device. (i) Iterative on-device computation cost during LLM FT: Existing device-server collaboration designs require the mobile devices to iteratively compute the activations across all training epochs. That induces enormous cumulative computation costs on the resource-constrained mobile device, which grows proportionally with LLM FT duration. (ii) Extra communication burden of activation transmissions: Both split learning and MobiLLM have to transfer full-dimension activations/gradients iteratively to the server, which incurs extra communication burdens. Especially when processing large batches with long sequence inputs, the communication cost may be too high to execute LLM FT with low-speed access or time-varying channels in wireless edge networks. (iii) Label or fine-tuned model exposure to the server: One major objective of on-device LLM FT is to keep data, labels, and fine-tuned models on the local mobile device. However, at the cost of server’s computing assistance, current device-server collaboration methods have the risk to disclose the true labels or fine-tuned models to the server. For instance, MobiLLM necessitates transmitting supervisory labels from the device to the server for parallel side-network training. Besides, the server can easily infer the fine-tuned model by adding up the tunable side network known by the server and the public frozen backbone model.
To address those issues above under device-server collaboration framework, in this paper, we propose PAE MobiLLM, a privacy-aware and efficient LLM FT approach, which can be deployed on the mobile device with the server’s help while maintaining data, labels and fine-tuned models on the local device [^1] Note that privacy awareness and protection in PAE MobiLLM means that it keeps data, labels, and fine-tuned models on the local device, which is similar to the data privacy protection definition in Google’s federated learning (FL) work  (19). Besides, the same as FL, PAE MobiLLM is obviously compatible with existing advanced privacy preserving approaches, such as differential privacy (DP) mechanisms  (10) (e.g., DP noises added on gradients, input, output or objective functions), secure multiparty computing (9), etc.. Our salient contributions are summarized as follows:
  • We develop an activation caching mechanism that allows the server to store intermediate activations from previously processed data samples. This not only accelerates the side-network training on the server side, but also enables the device to skip redundant forward propagation on the frozen backbone for recurring samples, reducing both computational and transmission costs.
  • We observe that a single activation dimension rather than the full activation matrix is good enough to guide adapter side-tuning of transformer-based models. Thus, we introduce a one-token activation transmission mechanism for shortcut connections between the mobile device and the server, which helps to significantly reduce communication costs and latency.
  • We design an additive adapter side-network, which can leverage the server’s computing capability to facilitate fine-tuning without disclosing ground-truth labels. The proposed additive adapter side-network consists of cascaded lightweight adapter modules that fuse model predictions via output difference-guided interpolation, which thereby eliminates the need for direct label transmission from the mobile device to the server.
  • We implement PAE MobiLLM on various mobile devices and development kits, including NVIDIA Jetson Xavier and Huawei MateBook, and evaluate its performance on sub-billion and billion-scale LLMs across diverse tasks and system settings. Experimental results demonstrate that compared with the second best LLM FT method on the mobile device, PAE MobiLLM achieves at least a 1.79×1.79$\times$1.79 × reduction in device-side computation, a 485.45×485.45$\times$485.45 × reduction in communication cost, and a 5.25×5.25$\times$5.25 × reduction in convergence time, while keeping data, true labels and fine-tuned models on the mobile device.
The remainder of the paper is organized as follows. Section II gives an introduction to LLM fine-tuning. Section III analyzes the limitations of existing works and presents our motivation. Section IV elaborates on our PAE MobiLLM design. Section V specifies experimental configurations. Section VI presents the performance evaluation. Section VII discusses our future work and Section VIII finally concludes the paper.
2. Preliminaries
2.1. Parameter- and Memory-Efficient FT
The memory footprint of LLM fine-tuning arises from three core components: model parameters, intermediate activations, and optimizer states. Extensive research has focused on optimizing these aspects, with Parameter-efficient Fine-tuning (PEFT) methods emerging as a prominent approach to reduce optimizer state memory and gradient computation costs by selectively updating sparse subsets of parameters while maintaining performance comparable to full fine-tuning. Techniques such as Adapter (12), which inserts bottleneck modules between transformer layers, LoRA (13), which decomposes weight updates using low-rank matrices, and BitFit (35), which exclusively trains bias terms, exemplify this paradigm. Further integration of model compression with PEFT—for instance, QLoRA (6)’s application of 4-bit quantization to the backbone network—has enabled significant reductions in parameter memory footprint. However, these approaches remain constrained by activation memory overhead as their trainable parameters remain embedded within the original model architecture, necessitating full backward passes through all layers. To address this limitation, recent innovations like ladder side-tuning (LST) (26) and LoSA (20) propose decoupling trainable components into a lightweight side network, creating a backpropagation highway that bypasses activation storage for the original model. Despite such advancements, practical deployment on resource-constrained edge devices faces unresolved challenges: low-bit quantization often demands specialized hardware support, and even memory-efficient designs like LST still exceed typical edge device capacities (16). These facts limit their feasibility in extremely resource-constrained environments.
Figure 1. Device-server collaborative LLM FT.
2.2. Device-Server Collaborative LLM FT
To enable device-side LLM adaptation while preserving private data, researchers are exploring on-device fine-tuning techniques. Given mobile hardware constraints, a common strategy involves partitioning the LLM and offloading part of the computation/storage burdens to adjacent servers or devices. Particularly, the U-shaped split learning (23) structure retains and updates only the few input and output transformer layers on the mobile device while delegating intermediate layer training to servers, ensuring that user data and labels remain local. As shown in Fig.  1(a), the training framework integrates U-shaped split learning with LoRA in a device-server collaborative setup, where activations of split layers are transmitted during forward propagation, and corresponding gradients are exchanged during backpropagation. When extended to multi-device scenarios, the LLM is further divided into layer-wise sub-models across devices, with forward/backward computations executed sequentially in a relay fashion (5). However, this layer-wise partitioning enforces strict sequential execution, and requires each device to update partial trainable parameters. In contrast, MobiLLM (16) introduces an innovative fine-tuning framework that decouples LLM fine-tuning into a frozen backbone model and a trainable side network, where the server hosts the adapter-based side network parallel to the device-side backbone (Fig. 1(b)) This framework maintains memory-efficient forward passes and preserves the original LLM functionality on resource-constrained devices, while offloading computation-intensive backpropagation to high-performance servers. Crucially, this parallel device-server training allows the device to execute lightweight inference through the frozen backbone while the server simultaneously computes gradients via the side network, which achieves efficient collaborative training without exposing raw input data.
2.3. “Pivot” Token in LLM architectures
Modern pre-trained LLMs predominantly adopt the Transformer architecture, where self-attention mechanisms capture dependencies across all tokens in an input sequence, enabling robust sentence-level comprehension. Fig. LABEL:fig:token_flow illustrates the activation dimension flow in a binary sentiment analysis task with a 3-token input sequence. After passing through the embedding layer, the input sequence is mapped to 3 tokens. The first transformer layer projects them to a shape of $[3,H]$, which remains unchanged in all subsequent transformer layers, where $H$ is the hidden size. At the final layer, a classification head projects the activations to $[3,2]$, requiring aggregation into a $[1,2]$ decision vector. Different architectures employ distinct token selector $S(\cdot)$ to extract global sentence representations, reflecting their pretraining objectives: For Auto-regressive models (e.g., GPT (2)) trained via left-to-right language modeling, they implicitly encode global context into the last token’s activation. The final token’s hidden state naturally accumulates bidirectional dependencies through stacked self-attention layers, serving as the de facto sentence representation. For auto-encoding models (e.g., BERT (7)) utilizing bidirectional context modeling, they prepend a dedicated [CLS] token to the input sequence. During pretraining, the [CLS] token’s embeddings are explicitly optimized to aggregate global sentence features through next-sentence prediction objectives, making it the canonical choice for downstream classification tasks. In conclusion, most LLMs designate a pivotal token ([CLS] in BERT, final token in GPT) to encode global context, making this token critical for the fine-tuning process compared to other positional dimensions.
Figure 2. Activation dimension changes in LLMs.
3. Motivation
While existing device-server collaboration methods enable LLM FT on the mobile device, there are still many critical issues for practical deployment in terms of computing efficiency, communications efficiency, and privacy awareness.
Observation 1: Iterative computation on the mobile device. LLM FT tasks typically demand dozens to hundreds of training epochs for convergence, requiring iterative training on the same set of local samples. While device-server collaborative frameworks offload partial computations to the server, mobile devices remain burdened by iterative local computation. For instance, in the split learning framework, the mobile device must recompute and transmit the intermidiate activations in every epoch—even for identical input samples—due to continuous parameter updates on the device-side model. MobiLLM eliminates device-side backpropagation by maintaining a frozen pretrained backbone, but it still enforces re-computation and transmission of static intermediate activations across training epochs by default. Table 1 compares communication and computational costs of device-only BitFit, split-learning BitFit, and MobiLLM. Since the required number of epochs in a fine-tuning session usually depends on dynamic factors like dataset size and hyperparameter configurations, prolonging on-device participation could make resource planning on memory/energy-constrained devices inherently unstable and inefficient.
Table 1. Communication and computational costs (On Xavier, task: MRPC@OPT-1.3B, batch size=8, seq len=256).
Method
Per-Batch
Comm. (MB)
Per-Epoch
Comp. (PFLOPS)
Total Comp.
BitFit-local - 18.4 208.3
BitFit-SL 33.4 1.5 17.4
MobilLM 200.4 9.5 124.5
Observation 2: High device-server communication burden. Current collaborative LLM FT methods require frequent device-server communications for exchanging intermediate activations or gradients—bidirectional in split learning (activations and gradients) and unidirectional in MobiLLM (activations only). This results in substantial transmission loads, especially for large models processing long sequences or large batches. For instance, in U-shaped split learning, fine-tuning OPT-1.3B with a batch size of 64 and token length of 512 requires transmitting up to 0.52 GB of data for just one batch. Despite advanced future wireless accessing technologies (e.g., 6G), communications remain a critical issue given the exponential growth of LLM parameters and edge deployment demands. However, a key architectural insight is often overlooked: each transformer layer inherently generates a “pivot” token that aggregates global semantic information and guides the model’s final decision. Existing methods transmit all token-level activations $[B,L,H]$ (Batch size $\times$ Seq Len $\times$ Hidden Size) by default, despite the pivot token’s compact representation $[B,H]$ being sufficient for LLM adaptation. That may offers a path to orders-of-magnitude efficiency gains - an opportunity yet unexplored in collaborative LLM FT paradigms.
Figure 3. An overview of the PAE MobiLLM system.
Observation 3: Unaware of label and fine-tuned model privacy. Existing device-server collaborative FT frameworks allow local retention of raw input data while requiring transmission of ground-truth labels to the server for loss calculation and partial parameter updates, e.g., updating deep layers in split learning or side-network parameters in MobiLLM. It may introduce systemic privacy vulnerabilities for label-sensitive applications, such as generative content creation or medical diagnostics. For instance, healthcare labels often contain highly sensitive information (e.g., disease diagnoses or genetic conditions). Moreover, these designs may inadvertently expose fine-tuned models. In MobiLLM, the server can combine the tuned side-network with the publicly known backbone model to replicate the device-specific fine-tuned LLM. Similarly, in split learning, although servers only update partial tunable parameters, they could maintain synchronized copies of client-side shallow layer initializations and mirror parameter updates to assemble the full model. While the U-shaped split learning variants (shown in Fig. 1(b)) keep labels locally, they fundamentally double the bidirectional activation/gradient transmission cycles, incurring excessive communication cost and increasing training latency as discussed in Observation 2.
4. PAE MobiLLM Design
4.1. PAE MobiLLM Overview
To address the observed issues, we propose PAE MobiLLM, a novel privacy-aware and efficient LLM FT design for the mobile device. PAE MobiLLM follows the device-server collaboration framework but effectively improves both computational and communication efficiency while ensuring data, label, i.e., $(\text{Input}_{x},\text{Label}_{y})$, and the fine-tuned model never leave the device. Briefly, as shown in Fig. 3, PAE MobiLLM assigns a frozen pre-trained model on the mobile device and lets it perform forward propagation only during the first epoch when batching over the local dataset. During this process, the mobile device forwards two types of information to the server: (1) the activation values ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) between transformer layers, which will be compressed by a token selector $S(\cdot)$ to reduce communication cost before transmissions; (2) the adjustment term $\Delta y$, as a linear calculation result of the backbone network’s forward inference result $y_{\text{pre}}$, the ground-truth label $\text{Label}_{y}$, and a locally generated random nonce $R$. PAE MobiLLM allocates all trainable modules to the server. During the first epoch, the server updates model parameters using ($\mathbf{A}_{1},\cdots,\mathbf{A}_{L},\Delta\mathbf{y}$) received from the device, while caching these ($\mathbf{A}_{1},\cdots,\mathbf{A}_{L},\Delta\mathbf{y}$) to construct a side-tuning guidance database. After the first epoch, the mobile device is freed, and the server will train the side-network iteratively until its convergence based on the cached activations. We will detail the key components, their functionalities and the procedure of PAE MobiLLM in the subsections below.
4.2. Key Computing-Efficient Components
4.2.1. Adapter Side-Tuning Network
Aiming to improve computing efficiency, PAE MobiLLM follows the device-server collaboration framework to offload the computing-intensive workload to the server. As shown in Fig.  3, PAE MobiLLM lets the mobile device only retain the frozen pre-trained model for forward inference and offloads the training and updating of the side-network to the server. In the side-tuning network, the trainable modules refine transformer-layer outputs from the backbone network, and enhance the representation quality for LLM FT. In PAE MobiLLM, for a pre-trained model with $L$ transformer layers, we construct $L$ adapter modules, where the $i$-th adapter module corresponds to the $i$-th transformer layer. The input and output dimensions of each adapter module match the hidden size of the transformer layers, ensuring seamless optimization of the backbone’s representations. The fine-tuning operations of the $i$-th layer is outlined in Fig.  4(a).
During the forward propagation process, let the $i$-th layer of the backbone network be denoted by $b_{i}$, and its hidden state output be represented by $h_{b_{i}}$. Similarly, the $i$-th layer of the side network adapter is denoted by $S_{i}$, and its output is represented by $h_{S_{i}}$. The input $S_{i\_in}$ to the $S_{i}$ is formulated as
(1) $S_{i\_in}=(1-\mu_{i})\cdot A_{i}+\mu_{i}\cdot h_{S_{i-1}},$
where $A_{i}$ is the activation value selected from the hidden state of the backbone $h_{b_{i}}$ by the token selector, $h_{S_{i-1}}$ is the output of the previous layer of the side network, and $\mu_{i}=\text{sigmoid}(\alpha_{i})$ is a gating parameter, where $\alpha_{i}$ denotes a learnable scalar initialized to zero.
Figure 4. Sketches of the adapter side-tuning.
Each adapter applies a down-projection matrix $\textbf{W}_{down}$$\in\mathbb{R}^{d*r}$ to reduce dimensionality, and then uses a non-linear activation function $\sigma(\cdot)$ and an up-projection matrix $\textbf{W}_{up}\in\mathbb{R}^{r*d}$ to restore the original dimension. Let $d$ be the hidden dimension of the input and $r$ denote the bottleneck dimension of the adapter. Here, $d$ is inherited from the pre-trained model, and $r$ is a tunable hyperparameter. The $i$-th adapter output includes a residual connection and is given by:
(2) $h_{S_{i}}=S_{i\_in}+\sigma(S_{i\_in}\cdot W_{\text{down}})W_{\text{up}}.$
After that, $h_{S_{i}}$ is passed to the next adapter layer, where it will be further trained until the final fine-tuned representation of the side network is obtained. Since all trainable modules are sequentially connected in the side network pathway, backpropagation will be completely executed within the server.
Therefore, under PAE MobiLLM collaborative network architecture, the mobile device only needs to take care of the forward pathway of the frozen backbone network based on the raw personal data, i.e., Inputx, thereby avoiding the high computational cost of tunable parts in LLM FT.
4.2.2. Activation Caching Mechanism on the Server
LLM FT typically requires multiple iterations of training, while iterative on-device training for activation updates will raise huge computational cost for the mobile device for either split learning or MobiLLM as discussed in Sec. 3. However, under the proposed PAE MobiLLM design, we observe that the backbone network and the side network work independently, and the backbone network on the mobile device remains frozen as shown in Fig. 3. This means that for the same data samples, activations remain unchanged across iterations.
Inspired by this observation, we propose an activation caching mechanism on the server. As shown in the rightmost part of Fig. 3, during the first epoch, the mobile device transmits intermediate activations ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) and the training target $\Delta\mathbf{y}$ to the server, while the server trains the side network and caches the received activations and $\Delta\mathbf{y}$. After the first epoch, since Inputx is the same and the backbone network is frozen on the mobile device, activations stay unchanged. So, the mobile device is freed. The server will train the side network iteratively until its convergence based on the cached ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$). The activation caching mechanism accelerates the side-tuning since the activation values from the backbone network are the same every iteration and the server can continuously conduct side-tuning using cached values without waiting for any activation updates.
Besides, due to activation caching, the mobile device only needs to compute the forward pass once and transmit the activation values to the server in the first epoch. That effectively improves computing efficiency and significantly reduces communication cost for the mobile device.
4.3. Key Communication-Efficient Components
Although device-server collaboration approaches enable LLM FT on the mobile device, they inevitably introduce extra communication costs due to the transmissions of activations. As the size of the model and the sequence length grows, the activation size increases as well and the activation transmissions may become a communication bottleneck, especially for low-speed access or time-varying channels in wireless edge networks. Aiming to minimize the communication cost, PAE MobiLLM uses a pivot token selector before activation transmissions.
The pivot token selector is originally used in pre-trained models for final decision-making. Depending on the model architecture, auto-regressive models (e.g., OPT) select the last token, while auto-encoding models (e.g., RoBERTa) use the [CLS] token. By extending those selection methods to transformer activations, PAE MobiLLM extracts the most representative token’s hidden states and transmits them to the server-side adapters, in order to reduce communication cost without sacrificing LLM FT performance. As shown in Fig.  4(b), for the activation tensors of shape $[B,L,H]$ (Batch size $\times$ Seq Len $\times$ Hidden Size), the token selector reduces the token dimension from $L$ to 1, and thus decreases the workload of activation transmissions from $B\times L\times H$ to $B\times H$. That generally reduces the communication cost by a factor of $1/L$ in PAE MobiLLM.
4.4. Privacy Awareness & Protection
Note that privacy awareness and protection in PAE MobiLLM means that it keeps data, labels and fine-tuned models on the local mobile device, which is similar to the data privacy protection defined in Google’s federated learning work (19). Besides, the same as federated learning, PAE MobiLLM is obviously compatible with existing advanced privacy preserving approaches, e.g., differential privacy mechanisms  (10), secure multiparty computing  (9), etc.
Following LST’s backbone and side-network design, MobiLLM has to transmit ground truth labels $(\text{Label}_{y})$ to the server and relies on the server to train the loss function $\text{Loss}(y_{\text{output}},\text{Label}_{y})$ as shown in Fig.  5(a). Although MobiLLM keeps the data on the mobile device by only transmitting the activations, the labels are disclosed to the server. Moreover, given the labels and public backbone network, it is very easy for the server who has trained side-network to infer the complete fine-tuned model. Thus, the privacy of label and fine-tuned model in MobiLLM cannot be protected.
Different from LST, Zhang et al. in (36) presented an additive learning based side-tuning. It decouples $y_{\text{output}}$ into $y_{\text{pre}}$ and $y_{\text{side}}$, i.e.,
(3) $y_{\text{output}}=y_{\text{pre}}+y_{\text{side}},$
where $y_{\text{output}}$ is the final model output, $y_{\text{pre}}$ is the original pre-trained model output, and $y_{\text{side}}$ is the additive trained output from the side network.
We follow this idea and apply it into PAE MobiLLM architecture. We let $y_{\text{pre}}$ be the output of backbone network, $y_{\text{side}}$ be the output of side network, and $y_{\text{side}}=y_{\text{output}}-y_{\text{pre}}$. Since the backbone network remains frozen on the mobile device, $y_{\text{pre}}$ is a constant during training in PAE MobiLLM. Thus, following $\text{Loss}(y_{\text{output}},\text{Label}_{y})$, the goal of the side network training is to minimize the loss function $\text{Loss}(y_{\text{side}},\text{Label}_{y}-y_{\text{pre}})$ as depicted in Fig.  3. Let $\Delta y$ denote the target tuning adjustment of the side network. Then, we have
(4) $\Delta y=\text{Label}_{y}-y_{\text{pre}}.$
If $\Delta y$ is sent to the server to guide side network training, $y_{\text{pre}}$ and $\text{Label}_{y}$ can stay on the mobile device. Then, the server may not be able to directly obtain $\text{Label}_{y}$.
However, the differential term $\Delta y$ may implicitly reveal certain trends due to the characteristics of the model’s probabilistic output. For example, in a three-class classification task, the predicted probability distribution is $y_{\text{pre}}=$ $[p_{1},p_{2},p_{3}]$ $=[0.3,0.6,0.1]$, where $p_{1}$ corresponds to Class 1, $p_{2}$ to Class 2, and $p_{3}$ to Class 3. If the value of $\text{Label}_{y}$ is $[0,1,0]$, then,
(5) $\Delta y=[0-p_{1},1-p_{2},0-p_{3}]=[-0.3,0.4,-0.1].$
Given the fact that $p_{1},p_{2},p_{3}$ are within the range $(0,1)$, $\Delta y$ in Eqn. (5) may leak the value of $\text{Label}_{y}$ due to the sign of $\Delta y$-values. Specifically, $(\text{Label}_{y}-y_{\text{pre}})$ always increases the probability of the true class while decreasing the probabilities of the other classes, which may expose the ground truth label.
Figure 5. How to execute on-device inferences in (a) MobiLLM; (b) PAE MobiLLM.
To address this issue, we introduce a random nonce $R$ generated by the mobile device, where each dimension is a real number randomly sampled from $(-1,1)$ (e.g., $[r_{1},r_{2},r_{3}]=[0.9,0.2,-0.3]$). Then, we let $\Delta y$ sent to the server be
(6) $\Delta y=\text{Label}_{y}-y_{\text{pre}}+R.$
Following Eqn. (6), in the previous example, $\Delta y$ becomes [0.6, 0.6, -0.4] and the server has no clue to infer the true label $\text{Label}_{y}$. After the server completes training and transmits the trained side network back to the mobile device, the final model output is computed as:
(7) $y_{\text{output}}=y_{\text{pre}}+y_{\text{side}}-R.$
[…]
4.5. PAE MobiLLM FT Procedure
Without loss of generality, we outline the LLM FT process of the proposed PAE MobiLLM design as follows:
(1) System Initialization (Device & Server). The device loads a pretrained LLM backbone compatible with local hardware constraints and generates a local privacy key $R$. The server then receives the device metadata to initialize a matched side network.
(2) First Epoch: Device-Server Collaborative Training.
(a) On-Device Forward Propagation. The device processes mini-batches through the frozen backbone, generating layer-wise activations $(\mathbf{A}_{1},\dots,\mathbf{A}_{L})$ and preliminary predictions $y_{\text{pre}}$ from the pretrained backbone model.
(b) Activation Processing & Transmission. The device applies activation selector $S(\cdot)$ to extract single-token activations from $h_{b_{i}}$rightarrow $A_{i}$ and computes delta target $\Delta y=\text{Label}_{y}-y_{\text{pre}}+R$. Then it transmits ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) to the server. Note that the device executes forward propagation asynchronously and progresses to the next batch without waiting for network transmission, avoiding blocking computation and reducing idle time during fine-tuning.
(c) Server-Side Side Network Training. The server merges the activations into the side network for forward propagation. By taking $\Delta y$ as the label, the server computes the training loss and performs backpropagation to update the side network parameters.
(d) Server-Side Activation Caching. The server stores activation-delta pairs ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) in the cache for subsequent training epochs.
(3) Subsequent Epoch: Server-Only Training. The server iterates on cached ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) pairs until convergence. This eliminates redundant device-side forward computation, thereby accelerating the training process.
(4) Side Network Deployment (Device & Server). The server returns final side network parameters to the device, which applies the privacy-preserving output fusion in Section 7 using the device-exclusive key $R$ (never exposed to the server). This ensures that the fine-tuned LLM is only available for the device while maintaining its compatibility with the original backbone model.
5. Experimental Setup
5.1. PAE MobiLLM Implementation
The experimental testbed consists of a server with NVIDIA A100 GPU (6912-core Ampere GPU with 40GB memory) and two representative mobile devices: (1) NVIDIA Jetson Xavier NX (6-core Carmel ARM CPU + 384-core Volta GPU, 8GB shared RAM) and (2) Huawei MateBook laptop (Intel Core i5-13500H CPU, 16GB RAM). Device-server communication employs Wi-Fi 5 (IEEE 802.11ac) with WebSocket protocol.
In particular, Jetson Xavier NX serves as the primary experimental platform in this work for evaluating the performance of GPU-accelerated LLM FT. Given the Jetson Xavier’s shared memory architecture, there is 4.6GB maximum GPU memory available for LLM training. The MateBook configuration evaluates CPU-only training scenarios, providing comparative insights into computation-constrained environments.
5.2. Models, Datasets and Parameters
Models: To systematically evaluate PAE MobiLLM’s performance, we employ two representative pre-trained architectures: the decoder-based OPT and encoder-based RoBERTa. Considering the Jetson Xavier platform’s 4.6GB GPU memory constraint, we select four model variants: OPT-1.3B, OPT-350M, RoBERTa-large (350M), and RoBERTa-base (125M), ensuring architectural diversity while maintaining mobile compatibility. All models are initialized via Huggingface’s Transformers library (31).
For the “pivot” token selection, we follow each architecture’s standard practice: OPT uses the last token’s hidden state (aligned with GPT series), while RoBERTa employs the [CLS] token embedding (consistent with BERT family) (17). That ensures architectural consistency across baseline comparisons.
Datasets: We take the GLUE benchmark (28) for the NLP task evaluation, which is widely used in the fine-tuning research for LLMs (37; 25; 26). It comprises seven classification tasks and one regression task, including linguistic acceptability (CoLA  (29)), sentiment analysis (SST-2 (24)), similarity and paraphrase (MRPC (8), QQP (14), STS-B (4)), and natural language inference (MNLI (30), QNLI (21), RTE (24)).
Parameters: To ensure fair comparison, all experiments share the same configurations unless specified: FP16 precision (except FP32 for CPU-only devices), 20 training epochs, batch size 8, learning rate 5e-4 (except 1e-6 for full FT), the maximum sequence length 256, and averaged 60 Mbps in-lab Wi-Fi transmission speed. We use NumPy’s random number generation library to generate $R$ with a fixed random seed of 0. Additionally, the internal low-rank hidden size of LoRA, MobiLLM, and PAE MobiLLM trainable modules is set to 64 by default.
(a) MRPC@Roberta-large
(b) MRPC@OPT-1.3B
(c) QNLI@Roberta-base
(d) QNLI@OPT-350M
Figure 6. Convergence performance comparisons over various models and tasks (On Xavier).
(a) MRPC@Roberta-large
(b) MRPC@OPT-1.3B
(c) QNLI@Roberta-base
(d) QNLI@OPT-350M
Figure 7. Convergence performance comparisons over various models and tasks (On laptop).
5.3. Baselines for Performance Comparison
We compare PAE MobiLLM with three baseline paradigms: i) Device only fine-tuning (denoted by ”-L”): Local execution using standard methods where all computation occurs on the device. ii) Split Learning (denoted by “-SL”): Integrates standard fine-tuning methods with U-shaped layer partitioning, retaining only the first and last transformer layers on devices while offloading intermediate layers to the server. iii) MobiLLM: Server-assisted collaborative fine-tuning leveraging a side network for LLM adaptation. The compared methods are implemented as follows:
Full-FT-L / Full-FT-SL: Full model fine-tuning.
LoRA-L/LoRA-SL (13): Inserts trainable low-rank matrices into the frozen backbone network.
BitFit-L/BitFit-SL (35): Updates only bias terms while freezing other weights.
MobiLLM-L (16): Trains both the backbone network and the side network locally on the mobile device.
MobiLLM (16): Server-assisted side-tuning which is evaluated without activation quantization for fair communication cost comparison.
6. Evaluation Results and Analysis
6.1. LLM FT Performance Analysis
We evaluate the training efficiency of PAE MobiLLM by measuring the total time required to reach target accuracy across diverse configurations involving two model architectures (OPT, RoBERTa), two tasks (MRPC, QNLI), and two mobile devices (NVIDIA Xavier, CPU-only laptop), as shown in Figs. 6 and 7. For clarity, we simplify the presentation by retaining only BitFit’s training curves (due to its similarity to LoRA) and excluding memory-intensive local FT baselines that exceed the mobile device’s memory capacity.
PAE MobiLLM is much faster than conventional split learning and server-assisted LLM FT baselines. Taking the FT of the MRPC task with the OPT-1.3B model (on Xavier) as an example, PAE MobiLLM is 15.5×15.5$\times$15.5 × faster than the second-fastest approach, BitFit-SL.Particularly, PAE MobiLLM demonstrates a two-phase acceleration pattern throughout the FT process. In the initial phase, PAE MobiLLM’s parallel scheduling design allows the mobile device to process subsequent data batches while the server simultaneously trains side networks using current-batch activations. This eliminates the idle time of the mobile device/server during activation transmissions and server computations, which is a delay bottleneck in traditional split learning frameworks. Besides, the one-token/pivot transmission in PAE MobiLLM remarkably reduces the transmission delay, which further speeds up the training. After the first epoch, PAE MobiLLM further strengthens its advantage through server-side activation caching. By reusing cached device-side activations, the server bypasses redundant computation and communications of activations for identical samples. This contrasts sharply with the baselines’ mandatory per-epoch device recomputations, where model updates force redundant activation recalculations even for static input data. By leveraging the server’s computational resources to handle cached activations, PAE MobiLLM progressively reduces the mobile device’s workload while maintaining stable convergence, ultimately achieving faster LLM FT compared to iterative recomputation and communications based approaches.
6.2. Advantages of Computing Efficiency
To evaluate the computational cost of the mobile device, we measure the total floating-point operations (FLOPs) required for LLM FT and use the MRPC task conducted over 20 epochs (Table 2) as a case study. The results demonstrate: (1) server-assisted methods significantly reduce on-device computation compared to fully local LLM FT; (2) PAE MobiLLM’s activation caching mechanism eliminates iterative on-device computations. Compared with MobiLLM, PAE MobiLLM achieves a 13.1×13.1$\times$13.1 × reduction in terms of on-device computing FLOPs, where the computing workload reduction is proportional to the number of training epochs. It indicates PAE MobiLLM’s one-time forward pass on-device computing design yields the highest computing efficiency for the mobile device across all LLM FT schemes.
Table 2. On-device computational cost to reach the target accuracy (On Xavier, task: MRPC).
Model OPT-1.3B RoBERTa-Large
Target Acc. 82.5% 87.1%
Method On-device Comp. (PFLOPS)
Per Epoch
Comp.
Total
Comp.
Per Epoch
Comp.
Total
Comp.
Full-FT-L 28.5 439.1 7.3 114.4
LoRA-L 20.5 235.5 5.2 57.6
BitFit-L 18.4 208.3 4.9 52.9
MobiLLM-L 9.7 127.3 2.5 33.6
Full-FT-SL 2.4 36.7 0.6 9.9
LoRA-SL 1.6 18.4 0.4 4.7
BitFit-SL 1.5 17.4 0.4 4.3
MobiLLM 9.5 124.5 2.4 32.5
PAE MobiLLM 9.5 9.5 2.4 2.4
6.3. Advantages of Communication Efficiency
As the results shown in Table 3, compared to device-server collaborative methods, PAE MobiLLM reduces at least 485.45×485.45$\times$485.45 × communication cost (PAE MobiLLM vs BitFit-SL, the OPT-1.3B model). We illustrate how PAE MobiLLM improves communication efficiency from the following three aspects.
Per-Batch Transmission Efficiency. Per-batch transmission measurements for OPT-1.3B (Table 3) reveal fundamental architectural differences. Split learning based methods necessitate dual activation-gradient exchanges and result in 33.4MB/batch communication load, while MobiLLM’s further escalates this burden due to its full-layer activation transmissions. In contrast, PAE MobiLLM leverages the pivot token selection to identify and only transmit the most informative activation dimension. As a result, it reduces per-batch traffic to under 1MB. That is a 97% reduction compared to split learning, and makes continuous device-server collaboration feasible even in low-speed or unstable wireless networks.
Sequence-Length Scalability. As illustrated in Fig. 10, conventional methods exhibit linear transmission growth with input sequence length (e.g., BitFit-SL increases 0.54MB per additional token). PAE MobiLLM’s dimension-consistent activation transfer mechanism breaks this dependency. That is, regardless of sequence length (64-512 tokens), communication remains stable below 1MB. Since split learning based approaches share the same transmission volume, we only take the BitFit-SL curve for illustrative purposes in Fig. 10. PAE MobiLLM’s fixed and low communication cost guarantees reliable provisioning for real-world scenarios with variable-length inputs, e.g., chatbots processing both short queries and long documents.
Total Communication Load. As demonstrated on the MRPC dataset (Table 3), all the other methods require iterative interactions with the server throughout the LLM FT process and lead to accumulated communication load proportional to sequence length $\times$ epoch number. By contrast, PAE MobiLLM processes the entire training dataset within the first epoch. It has no need for subsequent activation transmissions and only has a fixed initial cost. This promisingly enables precise estimation of total bandwidth requirements, which is critical for distributed training orchestration.
Table 3. Advantages in communication efficiency (On Xavier, PBC: Per Batch Comm. (MB), PEC: Per Epoch Comm. (GB), TC: Total Comm. (GB)).
Task MRPC@OPT-1.3B MRPC@OPT-350M MRPC@RoBERTa-Large MRPC@RoBERTa-Base
Target Acc. 82.5% 81.2% 87.1% 85.4%
Method Communication Load
PBC PEC TC PBC PEC TC PBC PEC TC PBC PEC TC
Full FT-SL 33.4 23.6 363.8 16.7 11.8 146.5 16.8 11.9 187.0 12.5 8.8 130.0
LoRA-SL 33.4 23.6 271.7 16.7 11.8 124.0 16.8 11.9 131.9 12.5 8.8 99.9
BitFit-SL 33.4 23.6 267.0 16.7 11.8 121.7 16.8 11.9 129.5 12.5 8.8 90.2
MobiLLM 200.4 141.7 1856.8 100.2 70.9 829.2 100.6 71.2 946.4 37.7 26.7 344.0
PAE MobiLLM 0.78 0.55 0.55 0.39 0.28 0.28 0.39 0.28 0.28 0.15 0.10 0.10
Figure 8. Impact of seq len on Comm. cost (On Xavier, OPT-1.3B).
Figure 9. Impact of seq len on execution time per batch (On Xavier, OPT-1.3B).
Figure 10. Impact of network speed on execution time per batch (On Xavier, OPT-1.3B).
6.4. Advantages of LLM FT Acceleration
[…]
Impact of Sequence Length. Figure 10 compares per-batch execution times across methods as input sequence length increases. Split learning based approaches (represented by BitFit-SL) exhibit linear growth in all three time components ($t_{1},t_{2},t_{3}$), which results in rapid total time escalation. MobiLLM’s communication burden also scales linearly with sequence length due to full activation transmissions, even though parallel computation partially mitigates the latency. PAE MobiLLM decouples the total execution time from sequence dependent communication cost, which constrains growth primarily to device-side forward computation. PAE MobiLLM achieves the lowest execution time across all tested lengths.
Table 4. Per-Batch Execution Time Breakdown.
Model Method Device: Xavier (Weak GPU) Device: Laptop (CPU-only) t3: Server Comp. (s) T:Total time (s)
t1: Comm.(s) t2: Comp. (s) t1: Comm. (s) t2: Comp. (s) Xavier Laptop
  OPT-1.3B LoRA-L - - - 27.51 - - 27.51
BitFit-L - - - 26.80 - - 26.80
MobiLLM-L - 4.32 - 19.29 - 4.32 19.29
Full-FT-SL 4.45 0.73 8.90 4.30 0.15 5.33 13.35
LoRA-SL 4.45 0.55 8.90 2.74 0.10 5.10 11.74
BitFit-SL 4.45 0.53 8.90 2.69 0.09 5.07 11.68
MobiLLM 26.72 2.85 53.44 10.30 0.01 26.72 53.44
PAE MobiLLM 0.10 2.86 0.20 10.34 0.01 2.86 10.34
  RoBERTa-large LoRA-L - 2.66 - 10.07 - 2.66 10.07
BitFit-L - 2.54 - 9.88 - 2.54 9.88
MobiLLM-L - 1.73 - 7.37 - 1.73 7.37
Full-FT-SL 2.22 0.30 4.44 1.25 0.06 2.58 5.75
LoRA-SL 2.22 0.21 4.44 0.58 0.06 2.49 5.08
BitFit-SL 2.22 0.19 4.44 0.53 0.04 2.45 5.01
MobiLLM 13.36 1.07 26.72 4.63 0.01 13.36 26.72
PAE MobiLLM 0.05 1.09 0.10 4.61 0.01 1.09 4.61
Impact of Wireless Transmission Speed. Figure 10 illustrates the per-batch execution time across methods under different wireless transmission speeds. Since the total execution time in BitFit-SL and MobiLLM is predominantly affected by the communication time, they suffer severe performance degradation under low-speed transmission conditions. For example, BitFit-SL shows an 8.75×8.75$\times$8.75 × slowdown at 10Mbps vs. 100Mbps. In contrast, PAE MobiLLM maintains stable total latency across different transmission speeds, demonstrating almost no variations from 10 Mbps to 100 Mbps. Such resilience stems from two facts: 1) single-token transmission takes very short time and thus minimizes the impact of wireless transmission speeds, and 2) parallelized scheduling overlaps communication phases with device-side computation, hiding transmission delays with on-device computation time that remains unaffected by transmission speed fluctuations.
7. Discussions and Future Work
The proposed PAE MobiLLM design demonstrates significant advancements in mobile LLM FT, but several practical considerations may merit deeper exploration to broaden its applicability.
Accuracy and Efficiency Trade-off: While PAE MobiLLM significantly improves the time, communication, and computational efficiency of on-device LLM fine-tuning, its lightweight side-network architecture and selective activation transmission strategy yield minor accuracy trade-offs (e.g., 0.4% and 1.2% reductions on QNLI tasks compared to SST2 and MobiLLM methods respectively). Our future work will explore adaptive mechanisms to balance efficiency and accuracy for precision-sensitive scenarios.
Server Storage: The server-side storage demands for caching activation data pose scalability challenges in resource constrained cloud environments. Techniques such as activation-aware quantization (e.g., mixed 2-4 bit precision per layer) and intelligent lifecycle management—where stale or low-impact activation pairs are automatically purged—could align storage costs with operational budgets without compromising adaptation quality.
Optimizing Side-Tuning: The server-side training pipeline also holds untapped optimization potential. Inspired by large-batch training strategies, dynamically scaling batch dimensions based on real-time GPU memory utilization could accelerate convergence while maintaining stability. For instance, progressively increasing batch sizes from 32 to 1024 following linear scaling rules may exploit hardware parallelism more effectively.
Mobile AI System Heterogeneity: Lastly, the growing heterogeneity of mobile hardware (e.g., varying CPU/TPU capabilities across devices) necessitates platform-agnostic computation strategies. Developing energy-proportional training protocols — where layer-wise computations adapt to device-specific power budgets via dynamic frequency scaling — could democratize efficient fine-tuning across diverse edge ecosystems. These directions collectively aim to transform theoretical advancements into deployable solutions for the evolving edge intelligence landscape.
8. Conclusion
This paper has presented PAE MobiLLM, a privacy-aware and efficient LLM FT on the mobile device via additive side-tuning. By decoupling the frozen backbone model on the mobile device from server-side side-network tuning, PAE MobiLLM offloads computation and memory-intensive workloads to the server while enabling practical LLM FT on mobile devices. To improve computing efficiency, PAE MobiLLM integrates activation caching on the server side to remove iterative on-device computations by reusing historical activations. To improve communication efficiency, PAE MobiLLM exploits one-token activation selection to compress activations and reduce transmission workloads. Aware of label and fine-tuned model privacy, PAE MobiLLM introduces the additive adapter side-network that isolates label-sensitive operations from the server. Experimental results show that PAE MobiLLM achieves at least a 1.79×1.79$\times$1.79 × reduction in device-side computation, a 485.45×485.45$\times$485.45 × reduction in communication cost, and a 5.25×5.25$\times$5.25 × acceleration in LLM FT convergence, while keeping data, training labels, and fine-tuned models on the mobile device. This work provides a deployable LLM FT solution to the resource-constrained mobile AI device whose computation/communication efficiency and data privacy are essential.

### v2

@@
 PAE MobiLLM: Privacy-Aware and Efficient LLM Fine-Tuning on the Mobile Device via Additive Side-Tuning
-Abstract.
+Abstract
 There is a huge gap between numerous intriguing applications fostered by on-device large language model (LLM) fine-tuning (FT) from fresh mobile data and the limited resources of a mobile device.
-While existing server-assisted methods (e.g., split learning or side-tuning) may enable LLM FT on the local mobile device, they suffer from heavy communication burdens of activation transmissions, and may disclose data, labels or fine-tuned models to the server.
+While existing server-assisted methods (e.g., split learning or side-tuning) may enable LLM FT on the local mobile device, they suffer from heavy communication burdens of activation transmissions, and may disclose data and labels to the server.
 To address those issues, we develop PAE MobiLLM, a privacy-aware and efficient LLM FT method which can be deployed on the mobile device via server-assisted additive side-tuning.
-The proposed PAE MobiLLM is featured by (i) notable communication and computing efficiency improvement; (ii) keeping data, true labels, and fine-tuned full models on the local mobile device; (iii) fast side-tuning on the server side.
-Briefly, PAE MobiLLM divides LLM fine-tune workloads between the mobile device and the server, i.e., the mobile device merely retains a frozen backbone model, while the computation-intensive trainable side-network is offloaded to the server.
 To further accelerate FT convergence and improve computing efficiency, PAE MobiLLM integrates activation caching on the server side, which allows the server to reuse historical activations and saves the mobile device from repeatedly computing forward passes for the recurring data samples.
-Besides, to reduce communication cost, PAE MobiLLM develops a one-token (i.e., “pivot” token) activation shortcut that transmits only a single activation dimension instead of full activation matrices to guide the side network tuning.
+Besides, to reduce communication cost, PAE MobiLLM develops an activation shortcut that transmits only the token involved in the loss calculation instead of full activation matrices to guide the side network tuning.
 Last but not least, PAE MobiLLM introduces the additive adapter side-network design which makes the server train the adapter modules based on device-defined prediction differences rather than raw ground-truth labels.
-In this way, the server can only assist device-defined side-network computing, and learn nothing about data, labels or fine-tuned models.
+In this way, the server can only assist device-defined side-network computing, and learn nothing about data and labels.
 Extensive experimental results demonstrate PAE MobiLLM’s superiority.
-LLM fine-tuning, Mobile device, Side-tuning, Computing and communication efficiency, Privacy awareness.
-Both authors contributed equally to this research.
-ccs: Computing methodologies Distributed artificial intelligence
-1. Introduction
-Large language models (LLMs) such as GPT (2), RoBERTa (17) and LLaMA (27) have recently demonstrated remarkable capabilities across various NLP tasks, including language representation (22), cross-task reasoning (34), and context-aware generation (18).
-This success stems from their two-stage adaptation paradigm: pretraining on large-scale public corpora followed by domain-specific fine-tuning (FT).
-While effective in centralized cloud environments, this conventional approach faces fundamental challenges in mobile computing scenarios, where user data is continuously generated across distributed mobile devices (e.g., smartphones and wearables), highly privacy-sensitive (e.g., health monitoring records and personal chat histories, and subject to stringent latency demands for intelligent services (15; 33).
-By adapting a local LLM with fresh data on the mobile device, on-device fine-tuning is emerging as both a practical necessity and a strategic shift toward personalized intelligence without compromising user privacy.
+Introduction
+Large language models (LLMs) like GPT (Brown et al.
+2020) and LLaMA (Touvron et al.
+2023) have recently demonstrated remarkable capabilities across various NLP tasks (Ren et al.
+2024; Ye 2024; Brown et al.
+2020) via pretraining on public corpora followed by domain-specific fine-tuning (FT).
+Although effective in centralized cloud environments, this paradigm faces fundamental challenges in mobile computing scenarios, where user data are continuously generated across distributed mobile devices (e.g., smartphones and wearables), highly privacy-sensitive (e.g., health monitoring records and personal chat histories), and subject to stringent latency demands for intelligent services (Li et al.
+2021; Ye et al.
+2020).
+How to enable LLM FT on the mobile device without compromising data privacy becomes an essential and popular research topic.
 Despite those promising applications, on-device LLM FT faces significant computing resource challenges.
@@
 Beyond storing the full model parameters, the predominant LLM training algorithm necessitates substantial memory for intermediate activations, gradients, and optimizer states.
-For example, fine-tuning a 7B-parameter LLM typically requires over 70 GB of GPU memory, far exceeding the capacity of most mobile devices, which often have 4–12 GB of RAM (32).
-A single training epoch for such an LLM could require hours of processing on mobile GPUs (3), which means full on-device fine-tuning cycles will take multiple days.
+For example, fine-tuning a 7B-parameter LLM typically requires over 70 GB of GPU memory, far exceeding the capacity of most mobile devices, which often have 4–12 GB of RAM (Xu et al.
+2024).
+A single training epoch for such an LLM could require hours of processing on mobile GPUs (Cai et al.
+2023), which means full on-device fine-tuning cycles will take multiple days.
 This memory usage and latency fundamentally conflict with practical deployment scenarios requiring continuous model adaptation.
 To enable LLM FT on a computing resource constrained mobile device, researchers are exploring to get some help from the high-performance computing server and collaborate the mobile device and the server via wireless communications supported by current mobile edge computing infrastructures.
-An intuitive solution is to integrate split learning (11) with PEFT, i.e., training shallow layers on the mobile device while offloading deeper layers to the server as shown in Fig.
-1(a).
+An intuitive solution is to integrate split learning (Gupta and Raskar 2018) with PEFT, i.e., training shallow layers on the mobile device while offloading deeper layers to the server.
 Yet, the sequential processing and bidirectional communications in this approach require fast wireless access and strict synchronization.
-Another device-server collaboration LLM FT approach is MobiLLM (16), which introduces server-assisted parallel side-tuning.
-Instead of splitting the model vertically by layers, MobiLLM decouples the trainable modules (hosted on the server) from the frozen backbone model (retained on the mobile device) as shown in Fig.
-1(b).
-During LLM FT, the mobile device executes forward passes on the backbone model and transmits intermediate activations to the server.
+Another device-server collaboration LLM FT approach is MobiLLM (Li et al.
+2025), which introduces server-assisted parallel side-tuning.
+Instead of splitting the model vertically by layers, MobiLLM decouples the trainable modules (hosted on the server) from the frozen backbone model (retained on the mobile device).
+During LLM FT, the mobile device executes forward propagation on the backbone model and transmits intermediate activations to the server.
 The server computes gradients and updates the side network via backpropagation, which helps the mobile device bypass the computing and memory intensive gradient calculations.
@@
 Especially when processing large batches with long sequence inputs, the communication cost may be too high to execute LLM FT with low-speed access or time-varying channels in wireless edge networks.
-(iii) Label or fine-tuned model exposure to the server: One major objective of on-device LLM FT is to keep data, labels, and fine-tuned models on the local mobile device.
-However, at the cost of server’s computing assistance, current device-server collaboration methods have the risk to disclose the true labels or fine-tuned models to the server.
-For instance, MobiLLM necessitates transmitting supervisory labels from the device to the server for parallel side-network training.
-Besides, the server can easily infer the fine-tuned model by adding up the tunable side network known by the server and the public frozen backbone model.
-To address those issues above under device-server collaboration framework, in this paper, we propose PAE MobiLLM, a privacy-aware and efficient LLM FT approach, which can be deployed on the mobile device with the server’s help while maintaining data, labels and fine-tuned models on the local device [^1] Note that privacy awareness and protection in PAE MobiLLM means that it keeps data, labels, and fine-tuned models on the local device, which is similar to the data privacy protection definition in Google’s federated learning (FL) work  (19).
-Besides, the same as FL, PAE MobiLLM is obviously compatible with existing advanced privacy preserving approaches, such as differential privacy (DP) mechanisms  (10) (e.g., DP noises added on gradients, input, output or objective functions), secure multiparty computing (9), etc..
-Our salient contributions are summarized as follows:
-  • We develop an activation caching mechanism that allows the server to store intermediate activations from previously processed data samples.
+(iii) Raw Label exposure to the server: One major objective of on-device LLM FT is to keep data and labels on the local mobile device.
+However, at the cost of server’s computing assistance, current device-server collaboration methods have the risk to disclose the true labels to the server.
+For instance, MobiLLM necessitates transmitting supervisory labels from the device to the server for side-network training.
+To tackle those issues, in this paper, we propose PAE MobiLLM, a privacy-aware and efficient LLM FT on the mobile device via device-server collaboration.
+PAE MobiLLM keeps data and labels local, while delegating memory/compute-intensive operations to an edge server via additive side-tuning.
+Our contributions are summarized as follows: (i) We design an additive adapter side-network, which can leverage the server’s computing capability to facilitate fine-tuning without disclosing ground-truth labels.
+The proposed additive adapter side-network consists of cascaded lightweight adapter modules that fuse model predictions via output difference-guided interpolation, which thereby eliminates the need for direct label transmission from the mobile device to the server.
+(ii) We introduce a token selection method that transmits only critical backbone-derived activations to supervise the side-network training at the server, thereby significantly reducing the communication workload.
+(iii) We develop an activation caching mechanism that allows the server to store intermediate activations from previously processed data samples.
 This not only accelerates the side-network training on the server side, but also enables the device to skip redundant forward propagation on the frozen backbone for recurring samples, reducing both computational and transmission costs.
-  • We observe that a single activation dimension rather than the full activation matrix is good enough to guide adapter side-tuning of transformer-based models.
-Thus, we introduce a one-token activation transmission mechanism for shortcut connections between the mobile device and the server, which helps to significantly reduce communication costs and latency.
-  • We design an additive adapter side-network, which can leverage the server’s computing capability to facilitate fine-tuning without disclosing ground-truth labels.
-The proposed additive adapter side-network consists of cascaded lightweight adapter modules that fuse model predictions via output difference-guided interpolation, which thereby eliminates the need for direct label transmission from the mobile device to the server.
-  • We implement PAE MobiLLM on various mobile devices and development kits, including NVIDIA Jetson Xavier and Huawei MateBook, and evaluate its performance on sub-billion and billion-scale LLMs across diverse tasks and system settings.
-Experimental results demonstrate that compared with the second best LLM FT method on the mobile device, PAE MobiLLM achieves at least a 1.79×1.79$\times$1.79 × reduction in device-side computation, a 485.45×485.45$\times$485.45 × reduction in communication cost, and a 5.25×5.25$\times$5.25 × reduction in convergence time, while keeping data, true labels and fine-tuned models on the mobile device.
-The remainder of the paper is organized as follows.
-Section II gives an introduction to LLM fine-tuning.
-Section III analyzes the limitations of existing works and presents our motivation.
-Section IV elaborates on our PAE MobiLLM design.
-Section V specifies experimental configurations.
-Section VI presents the performance evaluation.
-Section VII discusses our future work and Section VIII finally concludes the paper.
-2. Preliminaries
-2.1. Parameter- and Memory-Efficient FT
+(iv) We implement PAE MobiLLM on various mobile devices and evaluate its performance on sub-billion and billion-scale LLMs across diverse tasks and system settings.
+Experimental results demonstrate that compared to the second best LLM FT method on the mobile device, PAE MobiLLM achieves at least a $1.79\times$ reduction in device-side computation, a $17.06\times$ reduction in communication cost and a $5.25\times$ reduction in convergence time, while keeping data and true labels private.
+Preliminaries
+Parameter- and Memory-Efficient FT
 The memory footprint of LLM fine-tuning arises from three core components: model parameters, intermediate activations, and optimizer states.
 Extensive research has focused on optimizing these aspects, with Parameter-efficient Fine-tuning (PEFT) methods emerging as a prominent approach to reduce optimizer state memory and gradient computation costs by selectively updating sparse subsets of parameters while maintaining performance comparable to full fine-tuning.
-Techniques such as Adapter (12), which inserts bottleneck modules between transformer layers, LoRA (13), which decomposes weight updates using low-rank matrices, and BitFit (35), which exclusively trains bias terms, exemplify this paradigm.
-Further integration of model compression with PEFT—for instance, QLoRA (6)’s application of 4-bit quantization to the backbone network—has enabled significant reductions in parameter memory footprint.
+Techniques such as Adapter (Houlsby et al.
+2019), which inserts bottleneck modules between transformer layers, LoRA (Hu et al.
+2022), which decomposes weight updates using low-rank matrices, and BitFit (Zaken, Ravfogel, and Goldberg 2021), which exclusively trains bias terms, exemplify this paradigm.
 However, these approaches remain constrained by activation memory overhead as their trainable parameters remain embedded within the original model architecture, necessitating full backward passes through all layers.
-To address this limitation, recent innovations like ladder side-tuning (LST) (26) and LoSA (20) propose decoupling trainable components into a lightweight side network, creating a backpropagation highway that bypasses activation storage for the original model.
-Despite such advancements, practical deployment on resource-constrained edge devices faces unresolved challenges: low-bit quantization often demands specialized hardware support, and even memory-efficient designs like LST still exceed typical edge device capacities (16).
-These facts limit their feasibility in extremely resource-constrained environments.
-Figure 1. Device-server collaborative LLM FT.
-2.2. Device-Server Collaborative LLM FT
+To address this limitation, recent innovations like ladder side-tuning (LST) (Sung, Cho, and Bansal 2022) and LoSA (Mercea et al.
+2024) propose decoupling trainable components into a lightweight side network, creating a backpropagation highway that bypasses activation storage for the original model.
+Device-Server Collaborative LLM FT
 To enable device-side LLM adaptation while preserving private data, researchers are exploring on-device fine-tuning techniques.
 Given mobile hardware constraints, a common strategy involves partitioning the LLM and offloading part of the computation/storage burdens to adjacent servers or devices.
-Particularly, the U-shaped split learning (23) structure retains and updates only the few input and output transformer layers on the mobile device while delegating intermediate layer training to servers, ensuring that user data and labels remain local.
-As shown in Fig.
-1(a), the training framework integrates U-shaped split learning with LoRA in a device-server collaborative setup, where activations of split layers are transmitted during forward propagation, and corresponding gradients are exchanged during backpropagation.
-When extended to multi-device scenarios, the LLM is further divided into layer-wise sub-models across devices, with forward/backward computations executed sequentially in a relay fashion (5).
+Particularly, the U-shaped split learning (Ronneberger, Fischer, and Brox 2015) structure retains and updates only the few input and output transformer layers on the mobile device while delegating intermediate layer training to servers, ensuring that user data and labels remain local.
+When extended to multi-device scenarios, the LLM is further divided into layer-wise sub-models across devices, with forward/backward computations executed sequentially in a relay fashion (Chen et al.
+2023).
 However, this layer-wise partitioning enforces strict sequential execution, and requires each device to update partial trainable parameters.
-In contrast, MobiLLM (16) introduces an innovative fine-tuning framework that decouples LLM fine-tuning into a frozen backbone model and a trainable side network, where the server hosts the adapter-based side network parallel to the device-side backbone (Fig.
-1(b)) This framework maintains memory-efficient forward passes and preserves the original LLM functionality on resource-constrained devices, while offloading computation-intensive backpropagation to high-performance servers.
-Crucially, this parallel device-server training allows the device to execute lightweight inference through the frozen backbone while the server simultaneously computes gradients via the side network, which achieves efficient collaborative training without exposing raw input data.
-2.3. “Pivot” Token in LLM architectures
-Modern pre-trained LLMs predominantly adopt the Transformer architecture, where self-attention mechanisms capture dependencies across all tokens in an input sequence, enabling robust sentence-level comprehension.
-Fig.
-LABEL:fig:token_flow illustrates the activation dimension flow in a binary sentiment analysis task with a 3-token input sequence.
-After passing through the embedding layer, the input sequence is mapped to 3 tokens.
-The first transformer layer projects them to a shape of $[3,H]$, which remains unchanged in all subsequent transformer layers, where $H$ is the hidden size.
-At the final layer, a classification head projects the activations to $[3,2]$, requiring aggregation into a $[1,2]$ decision vector.
-Different architectures employ distinct token selector $S(\cdot)$ to extract global sentence representations, reflecting their pretraining objectives: For Auto-regressive models (e.g., GPT (2)) trained via left-to-right language modeling, they implicitly encode global context into the last token’s activation.
-The final token’s hidden state naturally accumulates bidirectional dependencies through stacked self-attention layers, serving as the de facto sentence representation.
-For auto-encoding models (e.g., BERT (7)) utilizing bidirectional context modeling, they prepend a dedicated [CLS] token to the input sequence.
-During pretraining, the [CLS] token’s embeddings are explicitly optimized to aggregate global sentence features through next-sentence prediction objectives, making it the canonical choice for downstream classification tasks.
-In conclusion, most LLMs designate a pivotal token ([CLS] in BERT, final token in GPT) to encode global context, making this token critical for the fine-tuning process compared to other positional dimensions.
-Figure 2. Activation dimension changes in LLMs.
-3. Motivation
+In contrast, MobiLLM (Li et al.
+2025) introduces an innovative framework that decouples LLM FT into a frozen backbone model and a trainable side network, where the server hosts the adapter-based side network parallel to the device-side backbone.
+This framework maintains memory-efficient forward passes and preserves the original LLM functionality on resource-constrained devices, while offloading computation-intensive backpropagation to high-performance servers.
+Crucially, this parallel device-server training allows the device to execute lightweight inference through the frozen backbone while the server simultaneously computes gradients via the side network, which achieves LLM FT without exposing raw input data.
+Figure 1: An overview of the PAE MobiLLM system.
+Motivation
 While existing device-server collaboration methods enable LLM FT on the mobile device, there are still many critical issues for practical deployment in terms of computing efficiency, communications efficiency, and privacy awareness.
@@
 MobiLLM eliminates device-side backpropagation by maintaining a frozen pretrained backbone, but it still enforces re-computation and transmission of static intermediate activations across training epochs by default.
-Table 1 compares communication and computational costs of device-only BitFit, split-learning BitFit, and MobiLLM.
-Since the required number of epochs in a fine-tuning session usually depends on dynamic factors like dataset size and hyperparameter configurations, prolonging on-device participation could make resource planning on memory/energy-constrained devices inherently unstable and inefficient.
-Table 1. Communication and computational costs (On Xavier, task: MRPC@OPT-1.3B, batch size=8, seq len=256).
-Method
-Per-Batch
-Comm. (MB)
-Per-Epoch
-Comp. (PFLOPS)
-Total Comp.
-BitFit-local - 18.4 208.3
-BitFit-SL 33.4 1.5 17.4
-MobilLM 200.4 9.5 124.5
 Observation 2: High device-server communication burden.
@@
 Despite advanced future wireless accessing technologies (e.g., 6G), communications remain a critical issue given the exponential growth of LLM parameters and edge deployment demands.
-However, a key architectural insight is often overlooked: each transformer layer inherently generates a “pivot” token that aggregates global semantic information and guides the model’s final decision.
-Existing methods transmit all token-level activations $[B,L,H]$ (Batch size $\times$ Seq Len $\times$ Hidden Size) by default, despite the pivot token’s compact representation $[B,H]$ being sufficient for LLM adaptation.
-That may offers a path to orders-of-magnitude efficiency gains - an opportunity yet unexplored in collaborative LLM FT paradigms.
-Figure 3. An overview of the PAE MobiLLM system.
-Observation 3: Unaware of label and fine-tuned model privacy.
+Observation 3: Unaware of label privacy.
 Existing device-server collaborative FT frameworks allow local retention of raw input data while requiring transmission of ground-truth labels to the server for loss calculation and partial parameter updates, e.g., updating deep layers in split learning or side-network parameters in MobiLLM.
 It may introduce systemic privacy vulnerabilities for label-sensitive applications, such as generative content creation or medical diagnostics.
-For instance, healthcare labels often contain highly sensitive information (e.g., disease diagnoses or genetic conditions).
-Moreover, these designs may inadvertently expose fine-tuned models.
-In MobiLLM, the server can combine the tuned side-network with the publicly known backbone model to replicate the device-specific fine-tuned LLM.
-Similarly, in split learning, although servers only update partial tunable parameters, they could maintain synchronized copies of client-side shallow layer initializations and mirror parameter updates to assemble the full model.
-While the U-shaped split learning variants (shown in Fig.
-1(b)) keep labels locally, they fundamentally double the bidirectional activation/gradient transmission cycles, incurring excessive communication cost and increasing training latency as discussed in Observation 2.
-4. PAE MobiLLM Design
-4.1. PAE MobiLLM Overview
+For instance, healthcare labels often contain highly sensitive information (e.g., disease diagnoses or genetic conditions), and in generation tasks, the labels can even be part of the raw input data.
+While the U-shaped split learning variants keep labels locally, they fundamentally double the bidirectional activation/gradient transmission cycles, incurring excessive communication cost and increasing training latency as discussed in Observation 2.
+PAE MobiLLM Design
+PAE MobiLLM Overview
 To address the observed issues, we propose PAE MobiLLM, a novel privacy-aware and efficient LLM FT design for the mobile device.
-PAE MobiLLM follows the device-server collaboration framework but effectively improves both computational and communication efficiency while ensuring data, label, i.e., $(\text{Input}_{x},\text{Label}_{y})$, and the fine-tuned model never leave the device.
+PAE MobiLLM follows the device-server collaboration framework but effectively improves both computational and communication efficiency while ensuring data, label, i.e., $(\text{Input}_{x},\text{Label}_{y})$ never leave the device.
 Briefly, as shown in Fig.
-3, PAE MobiLLM assigns a frozen pre-trained model on the mobile device and lets it perform forward propagation only during the first epoch when batching over the local dataset.
-During this process, the mobile device forwards two types of information to the server: (1) the activation values ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) between transformer layers, which will be compressed by a token selector $S(\cdot)$ to reduce communication cost before transmissions; (2) the adjustment term $\Delta y$, as a linear calculation result of the backbone network’s forward inference result $y_{\text{pre}}$, the ground-truth label $\text{Label}_{y}$, and a locally generated random nonce $R$.
+1, PAE MobiLLM assigns a frozen pre-trained model on the mobile device and lets it perform forward propagation only during the first epoch when batching over the local dataset.
+During this process, the mobile device forwards two types of information to the server: (1) the activation values ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) between transformer layers, which will be compressed by a token selector $S(\cdot)$ to reduce communication cost before transmissions; (2) the adjustment term $\Delta y$, as a linear calculation result of the backbone network’s forward inference result $y_{\text{pre}}$, the ground-truth label $\text{Label}_{y}$.
 PAE MobiLLM allocates all trainable modules to the server.
@@
 We will detail the key components, their functionalities and the procedure of PAE MobiLLM in the subsections below.
-4.2. Key Computing-Efficient Components
-4.2.1. Adapter Side-Tuning Network
+Key Computing-Efficient Components
+Adapter Side-Tuning Network
 Aiming to improve computing efficiency, PAE MobiLLM follows the device-server collaboration framework to offload the computing-intensive workload to the server.
 As shown in Fig.
-3, PAE MobiLLM lets the mobile device only retain the frozen pre-trained model for forward inference and offloads the training and updating of the side-network to the server.
-In the side-tuning network, the trainable modules refine transformer-layer outputs from the backbone network, and enhance the representation quality for LLM FT.
+1, PAE MobiLLM lets the mobile device only retain the frozen pre-trained model for forward inference and offloads the training and updating of the side-network to the server.
 In PAE MobiLLM, for a pre-trained model with $L$ transformer layers, we construct $L$ adapter modules, where the $i$-th adapter module corresponds to the $i$-th transformer layer.
@@
 The fine-tuning operations of the $i$-th layer is outlined in Fig.
-4(a).
+2.
 During the forward propagation process, let the $i$-th layer of the backbone network be denoted by $b_{i}$, and its hidden state output be represented by $h_{b_{i}}$.
@@
 The input $S_{i\_in}$ to the $S_{i}$ is formulated as
-(1) $S_{i\_in}=(1-\mu_{i})\cdot A_{i}+\mu_{i}\cdot h_{S_{i-1}},$
+$S_{i\_in}=(1-\mu_{i})\cdot A_{i}+\mu_{i}\cdot h_{S_{i-1}},$ (1)
 where $A_{i}$ is the activation value selected from the hidden state of the backbone $h_{b_{i}}$ by the token selector, $h_{S_{i-1}}$ is the output of the previous layer of the side network, and $\mu_{i}=\text{sigmoid}(\alpha_{i})$ is a gating parameter, where $\alpha_{i}$ denotes a learnable scalar initialized to zero.
-Figure 4. Sketches of the adapter side-tuning.
 Each adapter applies a down-projection matrix $\textbf{W}_{down}$$\in\mathbb{R}^{d*r}$ to reduce dimensionality, and then uses a non-linear activation function $\sigma(\cdot)$ and an up-projection matrix $\textbf{W}_{up}\in\mathbb{R}^{r*d}$ to restore the original dimension.
@@
 The $i$-th adapter output includes a residual connection and is given by:
-(2) $h_{S_{i}}=S_{i\_in}+\sigma(S_{i\_in}\cdot W_{\text{down}})W_{\text{up}}.$
+$h_{S_{i}}=S_{i\_in}+\sigma(S_{i\_in}\cdot W_{\text{down}})W_{\text{up}}.$ (2)
 After that, $h_{S_{i}}$ is passed to the next adapter layer, where it will be further trained until the final fine-tuned representation of the side network is obtained.
 Since all trainable modules are sequentially connected in the side network pathway, backpropagation will be completely executed within the server.
+Figure 2: Sketches of the adapter side-tuning.
 Therefore, under PAE MobiLLM collaborative network architecture, the mobile device only needs to take care of the forward pathway of the frozen backbone network based on the raw personal data, i.e., Inputx, thereby avoiding the high computational cost of tunable parts in LLM FT.
-4.2.2. Activation Caching Mechanism on the Server
-LLM FT typically requires multiple iterations of training, while iterative on-device training for activation updates will raise huge computational cost for the mobile device for either split learning or MobiLLM as discussed in Sec.
-3.
-However, under the proposed PAE MobiLLM design, we observe that the backbone network and the side network work independently, and the backbone network on the mobile device remains frozen as shown in Fig.
-3.
+Activation Caching Mechanism on the Server
+Under the proposed PAE MobiLLM design, we observe that the backbone network and the side network work independently, and the backbone network on the mobile device remains frozen as shown in Fig.
+1.
 This means that for the same data samples, activations remain unchanged across iterations.
@@
 As shown in the rightmost part of Fig.
-3, during the first epoch, the mobile device transmits intermediate activations ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) and the training target $\Delta\mathbf{y}$ to the server, while the server trains the side network and caches the received activations and $\Delta\mathbf{y}$.
+1, during the first epoch, the mobile device transmits intermediate activations ($\mathbf{A}_{1},\dots,\mathbf{A}_{L}$) and the training target $\Delta\mathbf{y}$ to the server, while the server trains the side network and caches the received activations and $\Delta\mathbf{y}$.
 After the first epoch, since Inputx is the same and the backbone network is frozen on the mobile device, activations stay unchanged.
@@
 The server will train the side network iteratively until its convergence based on the cached ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$).
-The activation caching mechanism accelerates the side-tuning since the activation values from the backbone network are the same every iteration and the server can continuously conduct side-tuning using cached values without waiting for any activation updates.
-Besides, due to activation caching, the mobile device only needs to compute the forward pass once and transmit the activation values to the server in the first epoch.
-That effectively improves computing efficiency and significantly reduces communication cost for the mobile device.
-4.3. Key Communication-Efficient Components
-Although device-server collaboration approaches enable LLM FT on the mobile device, they inevitably introduce extra communication costs due to the transmissions of activations.
-As the size of the model and the sequence length grows, the activation size increases as well and the activation transmissions may become a communication bottleneck, especially for low-speed access or time-varying channels in wireless edge networks.
-Aiming to minimize the communication cost, PAE MobiLLM uses a pivot token selector before activation transmissions.
-The pivot token selector is originally used in pre-trained models for final decision-making.
-Depending on the model architecture, auto-regressive models (e.g., OPT) select the last token, while auto-encoding models (e.g., RoBERTa) use the [CLS] token.
-By extending those selection methods to transformer activations, PAE MobiLLM extracts the most representative token’s hidden states and transmits them to the server-side adapters, in order to reduce communication cost without sacrificing LLM FT performance.
+Due to the activation caching mechanism, the mobile device only needs to compute the forward pass once and transmit the activation values to the server in the first epoch.
+That effectively improves computing efficiency and significantly reduces communication cost.
+Key Communication-Efficient Components
+Although device-server collaboration approaches enable LLM FT on the mobile device, they introduce additional communication costs due to the transmission of activations.
+Aiming to minimize the communication cost, PAE MobiLLM uses a token selector before activation transmissions.
+The token selector $S(\cdot)$ is originally used in pre-trained models for the alignment of the loss calculations.
 As shown in Fig.
-4(b), for the activation tensors of shape $[B,L,H]$ (Batch size $\times$ Seq Len $\times$ Hidden Size), the token selector reduces the token dimension from $L$ to 1, and thus decreases the workload of activation transmissions from $B\times L\times H$ to $B\times H$.
-That generally reduces the communication cost by a factor of $1/L$ in PAE MobiLLM.
-4.4. Privacy Awareness & Protection
-Note that privacy awareness and protection in PAE MobiLLM means that it keeps data, labels and fine-tuned models on the local mobile device, which is similar to the data privacy protection defined in Google’s federated learning work (19).
-Besides, the same as federated learning, PAE MobiLLM is obviously compatible with existing advanced privacy preserving approaches, e.g., differential privacy mechanisms  (10), secure multiparty computing  (9), etc.
+3, for FT training, the input sequence and the target label sequence often do not coincide in token counts.
+For the generation task, the input contains both the question and the answer concatenated as a complete semantic context, while the training target label sequence covers only the answer part.
+In classification tasks, each sentence corresponds to a single classification label, no matter how long it is.
+After forward propagation, to correctly calculate the supervised loss, it is necessary to align $\hat{y}$ with $\textit{Label}_{y}$ along the token dimension.
+Modern training procedures typically introduce a token-level selection function $S(\cdot)$ to extract a subset from $\hat{y}$ that matches the length of the label for calculating $\textit{Loss}(\hat{y},\textit{Label}_{y})$.
+We note that for the adapter module, applying the computation first and then selecting the tokens yields the same result as selecting the tokens before the computation.
+Based on this, we propose applying the token selector $S(\cdot)$, originally used in the output layer for loss computation, before transmission of intermediate activation values in each layer of the backbone network.
+By transmitting only the activations corresponding to token positions involved in loss computation, PAE MobiLLM significantly reduces communication cost without sacrificing LLM FT performance.
+Figure 3: Token-level selection for different LLM FT tasks.
+Privacy Awareness & Protection
 Following LST’s backbone and side-network design, MobiLLM has to transmit ground truth labels $(\text{Label}_{y})$ to the server and relies on the server to train the loss function $\text{Loss}(y_{\text{output}},\text{Label}_{y})$ as shown in Fig.
-5(a).
+4(a).
 Although MobiLLM keeps the data on the mobile device by only transmitting the activations, the labels are disclosed to the server.
-Moreover, given the labels and public backbone network, it is very easy for the server who has trained side-network to infer the complete fine-tuned model.
-Thus, the privacy of label and fine-tuned model in MobiLLM cannot be protected.
-Different from LST, Zhang et al. in (36) presented an additive learning based side-tuning. It decouples $y_{\text{output}}$ into $y_{\text{pre}}$ and $y_{\text{side}}$, i.e.,
-(3) $y_{\text{output}}=y_{\text{pre}}+y_{\text{side}},$
+Different from LST, Zhang et al. in (Zhang et al. 2020) presented an additive learning based side-tuning. It decouples $y_{\text{output}}$ into $y_{\text{pre}}$ and $y_{\text{side}}$, i.e.,
+$y_{\text{output}}=y_{\text{pre}}+y_{\text{side}},$ (3)
 where $y_{\text{output}}$ is the final model output, $y_{\text{pre}}$ is the original pre-trained model output, and $y_{\text{side}}$ is the additive trained output from the side network.
@@
 Thus, following $\text{Loss}(y_{\text{output}},\text{Label}_{y})$, the goal of the side network training is to minimize the loss function $\text{Loss}(y_{\text{side}},\text{Label}_{y}-y_{\text{pre}})$ as depicted in Fig.
-3.
+1.
 Let $\Delta y$ denote the target tuning adjustment of the side network.
 Then, we have
-(4) $\Delta y=\text{Label}_{y}-y_{\text{pre}}.$
+$\Delta y=\text{Label}_{y}-y_{\text{pre}}.$ (4)
 If $\Delta y$ is sent to the server to guide side network training, $y_{\text{pre}}$ and $\text{Label}_{y}$ can stay on the mobile device.
 Then, the server may not be able to directly obtain $\text{Label}_{y}$.
-However, the differential term $\Delta y$ may implicitly reveal certain trends due to the characteristics of the model’s probabilistic output.
-For example, in a three-class classification task, the predicted probability distribution is $y_{\text{pre}}=$ $[p_{1},p_{2},p_{3}]$ $=[0.3,0.6,0.1]$, where $p_{1}$ corresponds to Class 1, $p_{2}$ to Class 2, and $p_{3}$ to Class 3.
-If the value of $\text{Label}_{y}$ is $[0,1,0]$, then,
-(5) $\Delta y=[0-p_{1},1-p_{2},0-p_{3}]=[-0.3,0.4,-0.1].$
-Given the fact that $p_{1},p_{2},p_{3}$ are within the range $(0,1)$, $\Delta y$ in Eqn.
-(5) may leak the value of $\text{Label}_{y}$ due to the sign of $\Delta y$-values.
-Specifically, $(\text{Label}_{y}-y_{\text{pre}})$ always increases the probability of the true class while decreasing the probabilities of the other classes, which may expose the ground truth label.
-Figure 5. How to execute on-device inferences in (a) MobiLLM; (b) PAE MobiLLM.
-To address this issue, we introduce a random nonce $R$ generated by the mobile device, where each dimension is a real number randomly sampled from $(-1,1)$ (e.g., $[r_{1},r_{2},r_{3}]=[0.9,0.2,-0.3]$).
-Then, we let $\Delta y$ sent to the server be
-(6) $\Delta y=\text{Label}_{y}-y_{\text{pre}}+R.$
-Following Eqn.
-(6), in the previous example, $\Delta y$ becomes [0.6, 0.6, -0.4] and the server has no clue to infer the true label $\text{Label}_{y}$.
-After the server completes training and transmits the trained side network back to the mobile device, the final model output is computed as:
-(7) $y_{\text{output}}=y_{\text{pre}}+y_{\text{side}}-R.$
-[…]
-4.5. PAE MobiLLM FT Procedure
-Without loss of generality, we outline the LLM FT process of the proposed PAE MobiLLM design as follows:
-(1) System Initialization (Device & Server).
-The device loads a pretrained LLM backbone compatible with local hardware constraints and generates a local privacy key $R$.
-The server then receives the device metadata to initialize a matched side network.
-(2) First Epoch: Device-Server Collaborative Training.
-(a) On-Device Forward Propagation.
-The device processes mini-batches through the frozen backbone, generating layer-wise activations $(\mathbf{A}_{1},\dots,\mathbf{A}_{L})$ and preliminary predictions $y_{\text{pre}}$ from the pretrained backbone model.
-(b) Activation Processing & Transmission.
-The device applies activation selector $S(\cdot)$ to extract single-token activations from $h_{b_{i}}$rightarrow $A_{i}$ and computes delta target $\Delta y=\text{Label}_{y}-y_{\text{pre}}+R$.
-Then it transmits ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) to the server.
-Note that the device executes forward propagation asynchronously and progresses to the next batch without waiting for network transmission, avoiding blocking computation and reducing idle time during fine-tuning.
-(c) Server-Side Side Network Training.
-The server merges the activations into the side network for forward propagation.
-By taking $\Delta y$ as the label, the server computes the training loss and performs backpropagation to update the side network parameters.
-(d) Server-Side Activation Caching. The server stores activation-delta pairs ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) in the cache for subsequent training epochs.
-(3) Subsequent Epoch: Server-Only Training.
-The server iterates on cached ($\mathbf{A}_{1},\dots,\mathbf{A}_{L},\Delta\mathbf{y}$) pairs until convergence.
-This eliminates redundant device-side forward computation, thereby accelerating the training process.
-(4) Side Network Deployment (Device & Server).
-The server returns final side network parameters to the device, which applies the privacy-preserving output fusion in Section 7 using the device-exclusive key $R$ (never exposed to the server).
-This ensures that the fine-tuned LLM is only available for the device while maintaining its compatibility with the original backbone model.
-5. Experimental Setup
-5.1. PAE MobiLLM Implementation
-The experimental testbed consists of a server with NVIDIA A100 GPU (6912-core Ampere GPU with 40GB memory) and two representative mobile devices: (1) NVIDIA Jetson Xavier NX (6-core Carmel ARM CPU + 384-core Volta GPU, 8GB shared RAM) and (2) Huawei MateBook laptop (Intel Core i5-13500H CPU, 16GB RAM).
-Device-server communication employs Wi-Fi 5 (IEEE 802.11ac) with WebSocket protocol.
-In particular, Jetson Xavier NX serves as the primary experimental platform in this work for evaluating the performance of GPU-accelerated LLM FT.
-Given the Jetson Xavier’s shared memory architecture, there is 4.6GB maximum GPU memory available for LLM training.
-The MateBook configuration evaluates CPU-only training scenarios, providing comparative insights into computation-constrained environments.
-5.2. Models, Datasets and Parameters
+As shown in Fig.
+4(b), after the server completes training and transmits the trained side network back to the mobile device, the final model output is computed as Eqn.
+(3).
+Figure 4: How to execute on-device inferences in (a) MobiLLM; (b) PAE MobiLLM.
+Experimental Setup
+PAE MobiLLM Implementation
+The experimental testbed consists of a server with NVIDIA A100 GPU and two representative mobile devices: (1) NVIDIA Jetson Xavier NX (8GB RAM, 4.6GB GPU memory available for training) and (2) Huawei MateBook laptop (16GB RAM).
+In particular, Xavier NX serves as the primary platform for evaluating the performance of GPU-accelerated LLM FT.
+The MateBook represents CPU-only scenarios for comparison under computation-constrained environments.
+Task MRPC@OPT-1.3B Dialogsum@OPT-1.3B MRPC@RoBERTa-Large MRPC@RoBERTa-Base
+Target Acc. / ROUGE-1 82.5 (Acc.) 20.9 (ROUGE-1) 87.1 (Acc.) 85.4 (Acc.)
+Method Communication Load
+PBC PEC TC PBC PEC TC PBC PEC TC PBC PEC TC
+Full FT-SL 33.4 23.6 363.8 33.4 54.8 723.4 16.8 11.9 187.0 12.5 8.8 130.0
+LoRA-SL 33.4 23.6 271.7 33.4 54.8 695.9 16.8 11.9 131.9 12.5 8.8 99.9
+BitFit-SL 33.4 23.6 267.0 33.4 54.8 674.1 16.8 11.9 129.5 12.5 8.8 90.2
+MobiLLM 200.4 141.7 1856.8 200.4 329.2 3983.3 100.6 71.2 946.4 37.7 26.7 344.0
+PAE MobiLLM 0.78 0.55 0.55 24.1 39.5 39.5 0.39 0.28 0.28 0.15 0.10 0.10
+Table 1: Performance comparison of communication cost (Xavier, PBC: Per Batch Comm. (MB), PEC: Per Epoch Comm. (GB), TC: Total Comm. (GB)).
+Models, Datasets and Parameters
 Models: To systematically evaluate PAE MobiLLM’s performance, we employ two representative pre-trained architectures: the decoder-based OPT and encoder-based RoBERTa.
 Considering the Jetson Xavier platform’s 4.6GB GPU memory constraint, we select four model variants: OPT-1.3B, OPT-350M, RoBERTa-large (350M), and RoBERTa-base (125M), ensuring architectural diversity while maintaining mobile compatibility.
-All models are initialized via Huggingface’s Transformers library (31).
-For the “pivot” token selection, we follow each architecture’s standard practice: OPT uses the last token’s hidden state (aligned with GPT series), while RoBERTa employs the [CLS] token embedding (consistent with BERT family) (17).
-That ensures architectural consistency across baseline comparisons.
-Datasets: We take the GLUE benchmark (28) for the NLP task evaluation, which is widely used in the fine-tuning research for LLMs (37; 25; 26).
-It comprises seven classification tasks and one regression task, including linguistic acceptability (CoLA  (29)), sentiment analysis (SST-2 (24)), similarity and paraphrase (MRPC (8), QQP (14), STS-B (4)), and natural language inference (MNLI (30), QNLI (21), RTE (24)).
+All models are initialized via Huggingface’s Transformers library (Wolf et al.
+2019).
+Datasets: We take the GLUE benchmark (Wang et al.
+2018) and DialogSum dataset (Chen et al.
+2021) for the evaluation of NLP tasks, which are widely used in the fine-tuning research for LLM (Zhang et al.
+2023; Sun et al.
+2024; Sung, Cho, and Bansal 2022).
+GLUE benchmark comprises seven classification tasks and one regression task, including linguistic acceptability (CoLA  (Warstadt 2019)), sentiment analysis (SST-2 (Socher et al.
+2013)), similarity and paraphrase (MRPC (Dolan and Brockett 2005), QQP (Iyer et al.
+2017), STS-B (Cer et al.
+2017)), and natural language inference (MNLI (Williams, Nangia, and Bowman 2017), QNLI (Rajpurkar 2016), RTE (Socher et al.
+2013)).
+DialogSum includes summaries of real-word conversations on a diverse set of topics and scenarios to evaluate text-generation tasks.
 Parameters: To ensure fair comparison, all experiments share the same configurations unless specified: FP16 precision (except FP32 for CPU-only devices), 20 training epochs, batch size 8, learning rate 5e-4 (except 1e-6 for full FT), the maximum sequence length 256, and averaged 60 Mbps in-lab Wi-Fi transmission speed.
-We use NumPy’s random number generation library to generate $R$ with a fixed random seed of 0.
 Additionally, the internal low-rank hidden size of LoRA, MobiLLM, and PAE MobiLLM trainable modules is set to 64 by default.
-(a) MRPC@Roberta-large
-(b) MRPC@OPT-1.3B
-(c) QNLI@Roberta-base
-(d) QNLI@OPT-350M
-Figure 6. Convergence performance comparisons over various models and tasks (On Xavier).
-(a) MRPC@Roberta-large
-(b) MRPC@OPT-1.3B
-(c) QNLI@Roberta-base
-(d) QNLI@OPT-350M
-Figure 7. Convergence performance comparisons over various models and tasks (On laptop).
-5.3. Baselines for Performance Comparison
-We compare PAE MobiLLM with three baseline paradigms: i) Device only fine-tuning (denoted by ”-L”): Local execution using standard methods where all computation occurs on the device.
+Baselines for Performance Comparison
+We compare PAE MobiLLM with three baseline paradigms: i) Device only fine-tuning (denoted by ”-L”): Local execution using standard methods.
 ii) Split Learning (denoted by “-SL”): Integrates standard fine-tuning methods with U-shaped layer partitioning, retaining only the first and last transformer layers on devices while offloading intermediate layers to the server.
@@
 Full-FT-L / Full-FT-SL: Full model fine-tuning.
-LoRA-L/LoRA-SL (13): Inserts trainable low-rank matrices into the frozen backbone network.
-BitFit-L/BitFit-SL (35): Updates only bias terms while freezing other weights.
-MobiLLM-L (16): Trains both the backbone network and the side network locally on the mobile device.
-MobiLLM (16): Server-assisted side-tuning which is evaluated without activation quantization for fair communication cost comparison.
-6. Evaluation Results and Analysis
-6.1. LLM FT Performance Analysis
-We evaluate the training efficiency of PAE MobiLLM by measuring the total time required to reach target accuracy across diverse configurations involving two model architectures (OPT, RoBERTa), two tasks (MRPC, QNLI), and two mobile devices (NVIDIA Xavier, CPU-only laptop), as shown in Figs.
-6 and 7.
-For clarity, we simplify the presentation by retaining only BitFit’s training curves (due to its similarity to LoRA) and excluding memory-intensive local FT baselines that exceed the mobile device’s memory capacity.
-PAE MobiLLM is much faster than conventional split learning and server-assisted LLM FT baselines.
-Taking the FT of the MRPC task with the OPT-1.3B model (on Xavier) as an example, PAE MobiLLM is 15.5×15.5$\times$15.5 × faster than the second-fastest approach, BitFit-SL.Particularly, PAE MobiLLM demonstrates a two-phase acceleration pattern throughout the FT process.
-In the initial phase, PAE MobiLLM’s parallel scheduling design allows the mobile device to process subsequent data batches while the server simultaneously trains side networks using current-batch activations.
-This eliminates the idle time of the mobile device/server during activation transmissions and server computations, which is a delay bottleneck in traditional split learning frameworks.
-Besides, the one-token/pivot transmission in PAE MobiLLM remarkably reduces the transmission delay, which further speeds up the training.
-After the first epoch, PAE MobiLLM further strengthens its advantage through server-side activation caching.
-By reusing cached device-side activations, the server bypasses redundant computation and communications of activations for identical samples.
-This contrasts sharply with the baselines’ mandatory per-epoch device recomputations, where model updates force redundant activation recalculations even for static input data.
-By leveraging the server’s computational resources to handle cached activations, PAE MobiLLM progressively reduces the mobile device’s workload while maintaining stable convergence, ultimately achieving faster LLM FT compared to iterative recomputation and communications based approaches.
-6.2. Advantages of Computing Efficiency
-To evaluate the computational cost of the mobile device, we measure the total floating-point operations (FLOPs) required for LLM FT and use the MRPC task conducted over 20 epochs (Table 2) as a case study.
-The results demonstrate: (1) server-assisted methods significantly reduce on-device computation compared to fully local LLM FT; (2) PAE MobiLLM’s activation caching mechanism eliminates iterative on-device computations.
-Compared with MobiLLM, PAE MobiLLM achieves a 13.1×13.1$\times$13.1 × reduction in terms of on-device computing FLOPs, where the computing workload reduction is proportional to the number of training epochs.
-It indicates PAE MobiLLM’s one-time forward pass on-device computing design yields the highest computing efficiency for the mobile device across all LLM FT schemes.
-Table 2. On-device computational cost to reach the target accuracy (On Xavier, task: MRPC).
+LoRA-L/LoRA-SL (Hu et al. 2022): Inserts trainable low-rank matrices into the frozen backbone network.
+BitFit-L/BitFit-SL (Zaken, Ravfogel, and Goldberg 2021): Updates only bias terms while freezing other weights.
+MobiLLM-L (Li et al. 2025): Trains both the backbone network and the side network locally on the mobile device.
+MobiLLM (Li et al. 2025): Server-assisted side-tuning which is evaluated without activation quantization for fair communication cost comparison.
+Evaluation Results and Analysis
+Advantages of Computing Efficiency
 Model OPT-1.3B RoBERTa-Large
@@
 PAE MobiLLM 9.5 9.5 2.4 2.4
-6.3. Advantages of Communication Efficiency
-As the results shown in Table 3, compared to device-server collaborative methods, PAE MobiLLM reduces at least 485.45×485.45$\times$485.45 × communication cost (PAE MobiLLM vs BitFit-SL, the OPT-1.3B model).
-We illustrate how PAE MobiLLM improves communication efficiency from the following three aspects.
-Per-Batch Transmission Efficiency.
-Per-batch transmission measurements for OPT-1.3B (Table 3) reveal fundamental architectural differences.
-Split learning based methods necessitate dual activation-gradient exchanges and result in 33.4MB/batch communication load, while MobiLLM’s further escalates this burden due to its full-layer activation transmissions.
-In contrast, PAE MobiLLM leverages the pivot token selection to identify and only transmit the most informative activation dimension.
-As a result, it reduces per-batch traffic to under 1MB.
-That is a 97% reduction compared to split learning, and makes continuous device-server collaboration feasible even in low-speed or unstable wireless networks.
-Sequence-Length Scalability.
-As illustrated in Fig.
-10, conventional methods exhibit linear transmission growth with input sequence length (e.g., BitFit-SL increases 0.54MB per additional token).
-PAE MobiLLM’s dimension-consistent activation transfer mechanism breaks this dependency.
-That is, regardless of sequence length (64-512 tokens), communication remains stable below 1MB.
-Since split learning based approaches share the same transmission volume, we only take the BitFit-SL curve for illustrative purposes in Fig.
-10.
-PAE MobiLLM’s fixed and low communication cost guarantees reliable provisioning for real-world scenarios with variable-length inputs, e.g., chatbots processing both short queries and long documents.
-Total Communication Load.
-As demonstrated on the MRPC dataset (Table 3), all the other methods require iterative interactions with the server throughout the LLM FT process and lead to accumulated communication load proportional to sequence length $\times$ epoch number.
-By contrast, PAE MobiLLM processes the entire training dataset within the first epoch.
-It has no need for subsequent activation transmissions and only has a fixed initial cost.
-This promisingly enables precise estimation of total bandwidth requirements, which is critical for distributed training orchestration.
-Table 3. Advantages in communication efficiency (On Xavier, PBC: Per Batch Comm. (MB), PEC: Per Epoch Comm. (GB), TC: Total Comm. (GB)).
-Task MRPC@OPT-1.3B MRPC@OPT-350M MRPC@RoBERTa-Large MRPC@RoBERTa-Base
-Target Acc. 82.5% 81.2% 87.1% 85.4%
-Method Communication Load
-PBC PEC TC PBC PEC TC PBC PEC TC PBC PEC TC
-Full FT-SL 33.4 23.6 363.8 16.7 11.8 146.5 16.8 11.9 187.0 12.5 8.8 130.0
-LoRA-SL 33.4 23.6 271.7 16.7 11.8 124.0 16.8 11.9 131.9 12.5 8.8 99.9
-BitFit-SL 33.4 23.6 267.0 16.7 11.8 121.7 16.8 11.9 129.5 12.5 8.8 90.2
-MobiLLM 200.4 141.7 1856.8 100.2 70.9 829.2 100.6 71.2 946.4 37.7 26.7 344.0
-PAE MobiLLM 0.78 0.55 0.55 0.39 0.28 0.28 0.39 0.28 0.28 0.15 0.10 0.10
-Figure 8. Impact of seq len on Comm. cost (On Xavier, OPT-1.3B).
-Figure 9. Impact of seq len on execution time per batch (On Xavier, OPT-1.3B).
-Figure 10. Impact of network speed on execution time per batch (On Xavier, OPT-1.3B).
-6.4. Advantages of LLM FT Acceleration
-[…]
-Impact of Sequence Length.
-Figure 10 compares per-batch execution times across methods as input sequence length increases.
-Split learning based approaches (represented by BitFit-SL) exhibit linear growth in all three time components ($t_{1},t_{2},t_{3}$), which results in rapid total time escalation.
-MobiLLM’s communication burden also scales linearly with sequence length due to full activation transmissions, even though parallel computation partially mitigates the latency.
-PAE MobiLLM decouples the total execution time from sequence dependent communication cost, which constrains growth primarily to device-side forward computation.
-PAE MobiLLM achieves the lowest execution time across all tested lengths.
-Table 4. Per-Batch Execution Time Breakdown.
-Model Method Device: Xavier (Weak GPU) Device: Laptop (CPU-only) t3: Server Comp. (s) T:Total time (s)
+Table 2: On-device computational cost (Xavier, task: MRPC).
+To evaluate the computational cost of the mobile device, we measure the total floating-point operations (FLOPs) required for LLM FT and use the MRPC task conducted over 20 epochs (Table 2) as a case study.
+The results demonstrate: (1) server-assisted methods significantly reduce on-device computation compared to fully local LLM FT; (2) PAE MobiLLM’s activation caching mechanism eliminates iterative on-device computations.
+Compared with MobiLLM, PAE MobiLLM achieves a $13.1\times$ reduction in terms of on-device computing FLOPs, where the computing workload reduction is proportional to the number of training epochs.
+It indicates PAE MobiLLM’s one-time forward pass on-device computing design yields the highest computing efficiency for the mobile device across all LLM FT schemes.
+Advantages of Communication Efficiency
+As the results shown in Table 1, compared to device-server collaborative methods, PAE MobiLLM reduces at least $17.06\times$ communication cost (PAE MobiLLM vs BitFit-SL, the OPT-1.3B model, and Dialogsum task).
+We illustrate how PAE MobiLLM improves communication efficiency from the following two aspects.
+MRPC@OPT-1.3B (Xavier)
+(a) *
+MRPC@OPT-1.3B (laptop)
+(b) *
+QNLI@RoBERTa-Base (Xavier)
+(c) *
+QNLI@RoBERTa-Base (laptop)
+(d) *
+Figure 5: Convergence performance comparisons over various models, tasks, and devices.
+Method Device: Xavier (Weak GPU) Device: Laptop (CPU-only)
+t3: Server
+Comp. (s)
+T: Total time (s)
 t1: Comm.(s) t2: Comp. (s) t1: Comm.(s) t2: Comp. (s) Xavier Laptop
-  OPT-1.3B LoRA-L - - - 27.51 - - 27.51
+LoRA-L - - - 27.51 - - 27.51
 BitFit-L - - - 26.80 - - 26.80
@@
 PAE MobiLLM 0.10 2.86 0.20 10.34 0.01 2.86 10.34
-  RoBERTa-large LoRA-L - 2.66 - 10.07 - 2.66 10.07
-BitFit-L - 2.54 - 9.88 - 2.54 9.88
-MobiLLM-L - 1.73 - 7.37 - 1.73 7.37
-Full-FT-SL 2.22 0.30 4.44 1.25 0.06 2.58 5.75
-LoRA-SL 2.22 0.21 4.44 0.58 0.06 2.49 5.08
-BitFit-SL 2.22 0.19 4.44 0.53 0.04 2.45 5.01
-MobiLLM 13.36 1.07 26.72 4.63 0.01 13.36 26.72
-PAE MobiLLM 0.05 1.09 0.10 4.61 0.01 1.09 4.61
-Impact of Wireless Transmission Speed.
-Figure 10 illustrates the per-batch execution time across methods under different wireless transmission speeds.
-Since the total execution time in BitFit-SL and MobiLLM is predominantly affected by the communication time, they suffer severe performance degradation under low-speed transmission conditions.
-For example, BitFit-SL shows an 8.75×8.75$\times$8.75 × slowdown at 10Mbps vs.
-100Mbps.
-In contrast, PAE MobiLLM maintains stable total latency across different transmission speeds, demonstrating almost no variations from 10 Mbps to 100 Mbps.
-Such resilience stems from two facts: 1) single-token transmission takes very short time and thus minimizes the impact of wireless transmission speeds, and 2) parallelized scheduling overlaps communication phases with device-side computation, hiding transmission delays with on-device computation time that remains unaffected by transmission speed fluctuations.
-7. Discussions and Future Work
-The proposed PAE MobiLLM design demonstrates significant advancements in mobile LLM FT, but several practical considerations may merit deeper exploration to broaden its applicability.
-Accuracy and Efficiency Trade-off: While PAE MobiLLM significantly improves the time, communication, and computational efficiency of on-device LLM fine-tuning, its lightweight side-network architecture and selective activation transmission strategy yield minor accuracy trade-offs (e.g., 0.4% and 1.2% reductions on QNLI tasks compared to SST2 and MobiLLM methods respectively).
-Our future work will explore adaptive mechanisms to balance efficiency and accuracy for precision-sensitive scenarios.
-Server Storage: The server-side storage demands for caching activation data pose scalability challenges in resource constrained cloud environments.
-Techniques such as activation-aware quantization (e.g., mixed 2-4 bit precision per layer) and intelligent lifecycle management—where stale or low-impact activation pairs are automatically purged—could align storage costs with operational budgets without compromising adaptation quality.
-Optimizing Side-Tuning: The server-side training pipeline also holds untapped optimization potential.
-Inspired by large-batch training strategies, dynamically scaling batch dimensions based on real-time GPU memory utilization could accelerate convergence while maintaining stability.
-For instance, progressively increasing batch sizes from 32 to 1024 following linear scaling rules may exploit hardware parallelism more effectively.
-Mobile AI System Heterogeneity: Lastly, the growing heterogeneity of mobile hardware (e.g., varying CPU/TPU capabilities across devices) necessitates platform-agnostic computation strategies.
-Developing energy-proportional training protocols — where layer-wise computations adapt to device-specific power budgets via dynamic frequency scaling — could democratize efficient fine-tuning across diverse edge ecosystems.
-These directions collectively aim to transform theoretical advancements into deployable solutions for the evolving edge intelligence landscape.
-8. Conclusion
+Table 3: Per-Batch Execution Time Breakdown (task:SST-2@OPT-1.3B).
+Per-Batch Transmission Efficiency.
+Per-batch transmission measurements for OPT-1.3B (Table 1) reveal fundamental architectural differences.
+Split learning based methods necessitate dual activation-gradient exchanges and result in 33.4MB/batch communication load, while MobiLLM’s further escalates this burden due to its full-layer activation transmissions.
+In contrast, PAE MobiLLM leverages the token selector to identify and transmit only the activation dimensions involved in the loss computation.
+Compared to split learning, it reduces per-batch traffic by at least 27.8% in generation task and even up to 97% in classification tasks, which makes such device-server collaboration based LLM FT feasible in low-speed or unstable wireless networks.
+Total Communication Load.
+As demonstrated in the Table 1, all the other methods require iterative interactions with the server throughout the LLM FT process and lead to accumulated communication load proportional to the number of epoch.
+By contrast, PAE MobiLLM processes the entire training dataset within the first epoch.
+It has no need for subsequent activation transmissions and only has a fixed initial cost.
+This promisingly enables precise estimation of total bandwidth requirements, which is critical for distributed training orchestration.
+Advantages of LLM FT Acceleration
+We evaluate the training efficiency of PAE MobiLLM by measuring the total time required to reach target accuracy across diverse configurations involving two model architectures (OPT, RoBERTa), two tasks (MRPC, QNLI), and two mobile devices (NVIDIA Xavier, CPU-only laptop), as shown in Fig.
+5.
+For clarity, we simplify the presentation by retaining only BitFit’s training curves (due to its similarity to LoRA) and excluding memory-intensive local FT baselines that exceed the mobile device’s memory capacity.
+PAE MobiLLM is much faster than conventional split learning and server-assisted LLM FT baselines.
+Taking the FT of the MRPC task with the OPT-1.3B model (on Xavier) as an example, PAE MobiLLM is $15.5\times$ faster than the second-fastest approach, BitFit-SL.
+Particularly, PAE MobiLLM demonstrates a two-phase acceleration pattern throughout the FT process.
+In the initial phase, PAE MobiLLM’s parallel scheduling design allows the mobile device to process subsequent data batches while the server simultaneously trains side networks using current-batch activations.
+This eliminates the idle time of the mobile device/server during activation transmissions and server computations, which is a delay bottleneck in traditional split learning frameworks.
+Besides, using token selector before transmission in PAE MobiLLM remarkably reduces the transmission delay, which further speeds up the training.
+After the first epoch, PAE MobiLLM leverages server-side caching to bypass redundant device-side computation and communication for identical sample activations.
+By leveraging the server’s computational resources to perform independent iterative training, PAE MobiLLM achieves faster LLM FT while maintaining stable convergence, which outperforms iterative recomputation and communication based LLM FT approaches.
+Breakdown of Per-Batch Execution Time.
+Table 3 compares the time costs, including communication time $t_{1}$, on-device computation time $t_{2}$, and server-side processing time $t_{3}$, for forward and backward passes across different approaches.
+Split learning based methods execute all the stages sequentially and result in cumulative latency of the three parts.
+While MobiLLM supports device-server parallel computation, its heavy activation transmission burden dominates its total LLM FT latency, especially when processing long sequences or large batches.
+Unlike those designs, PAE MobiLLM alleviates this bottleneck through token selection, reduces per-batch data transmission and makes transmission latency negligible.
+For the on-device computing time, both MobiLLM and PAE MobiLLM remove backward propagation workloads from the mobile devices, which only need to conduct forward inference computations.
+That helps to simplify on-device operations, making MobiLLM and PAE MobiLLM naturally compatible with emerging inference accelerators tailored for mobile devices (e.g., mobile TPUs).
+As shown in Table 3, PAE MobiLLM consistently achieves the shortest per-batch execution time across all tested devices and model scales, demonstrating superior training efficiency.
+LLM FT Performance Analysis
+PAE MobiLLM introduces three key components beyond the MobiLLM framework, i.e., the additive side network, the token selector and the server cache, respectively.
+We conducted ablation studies to assess the contribution of each component to LLM FT performance.
+For the GLUE benchmark, we report the average score, which combines accuracy on MNLI, QQP, QNLI, SST-2, MRPC, and RTE, Pearson correlation on STS-B, and Matthews correlation on CoLA.
+For the DialogSum dataset, we use ROUGE scores (R1/R2/RL) as the accuracy metric.
+As shown in Table 4, both the server cache and the token selector significantly improve training efficiency without degrading LLM FT performance.
+The additive side network also consistently maintains competitive performance across different datasets.
+The experimental results demonstrate that the proposed PAE MobiLLM achieves excellent training efficiency and enhanced data privacy without sacrificing LLM FT performance.
+Model OPT-1.3B RoBERTa-Large
+Data
+GLUE
+Dialogsum
+(R1/R2/RL)
+GLUE
+PAE MobiLLM 81.1 21.1 / 7.9 / 17.1 86.8
+w/o. ASN 81.4 21.3 / 7.5 / 17.6 86.9
+w/o. TS 81.2 21.1 / 7.7 / 17.0 86.7
+w/o. SC 81.1 21.0 / 7.7 / 17.1 86.8
+w/o. ASN/ TS/ SC 81.4 21.3 / 7.6 / 17.5 86.6
+Table 4: Ablation studies of the proposed PAE MobiLLM designs (ASN: Additive Side Network, TS: Token Selector, SC: Sever Cache, w/o. : Without).
+Conclusion
 This paper has presented PAE MobiLLM, a privacy-aware and efficient LLM FT on the mobile device via additive side-tuning.
@@
 To improve computing efficiency, PAE MobiLLM integrates activation caching on the server side to remove iterative on-device computations by reusing historical activations.
-To improve communication efficiency, PAE MobiLLM exploits one-token activation selection to compress activations and reduce transmission workloads.
-Aware of label and fine-tuned model privacy, PAE MobiLLM introduces the additive adapter side-network that isolates label-sensitive operations from the server.
-Experimental results show that PAE MobiLLM achieves at least a 1.79×1.79$\times$1.79 × reduction in device-side computation, a 485.45×485.45$\times$485.45 × reduction in communication cost, and a 5.25×5.25$\times$5.25 × acceleration in LLM FT convergence, while keeping data, training labels, and fine-tuned models on the mobile device.
-This work provides a deployable LLM FT solution to the resource-constrained mobile AI device whose computation/communication efficiency and data privacy are essential.
+To improve communication efficiency, PAE MobiLLM exploits token selection to compress activations and reduce transmission workloads.
+Aware of label privacy, PAE MobiLLM introduces the additive adapter side-network that isolates label-sensitive operations from the server.
+Experimental results show that PAE MobiLLM achieves at least a $1.79\times$ reduction in device-side computation, a $17.06\times$ reduction in communication cost, and a $5.25\times$ acceleration in LLM FT convergence, while keeping data and labels private.